### v1

# Scaling Laws for Online Advertisement Retrieval

Abstract.
The scaling law is a notable property of neural network models that researchers have discovered, which greatly propels the development of large language models. Scaling laws hold great promise in guiding model design and resource allocation. Recent research increasingly shows that scaling laws are not limited to NLP tasks or Transformer architectures; they also apply to domains such as recommendation, where there are scaling laws relating to the amount of training data, model size, and model performance metrics. However, there is still a lack of literature on scaling law research in industrial online advertisement retrieval systems. This may be because 1) identifying the scaling law for resource cost and online revenue is often expensive in both time and training resources for large-scale industrial applications, and 2) varying settings for different systems prevent the scaling law from being applied across various scenarios. To address these issues, we propose a lightweight paradigm to identify the scaling law of online revenue and machine cost for a certain online advertisement retrieval scenario with a low experimental cost. Specifically, we focus on a sole factor (FLOPs) and propose an offline metric named $R/R^{*}$ that exhibits a high linear correlation with online revenue for retrieval models. We estimate the machine cost offline via a simulation algorithm. Thus, we can transform most online experiments into low-cost offline experiments. We conduct comprehensive experiments to verify the effectiveness of our proposed metric $R/R^{*}$ and to identify the scaling law in the online advertisement retrieval system of Kuaishou. With the scaling law, we demonstrate practical applications for ROI-constrained model designing and multi-scenario resource allocation in the Kuaishou advertising system. To the best of our knowledge, this is the first work to study the scaling laws for online advertisement retrieval of real-world large-scale systems, showing great potential for scaling law in advertising system optimization.
Scaling Law, Online Advertisement Retrieval, Resource Allocation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning
1. Introduction
The neural scaling laws, describing how neural network performance changes with key factors (e.g. model size, dataset size, computational cost), have been discovered in various research areas (Kaplan et al., 2020; Hoffmann et al., 2022; Shin et al., 2023; Isik et al., 2024; Zhang et al., 2024a; Fang et al., 2024). Early research (Hestness et al., 2017) shows that the neural network performance is predictable when scaling training data size in various tasks such as neural machine translation and language modeling. Kaplan et al.((2020)(Kaplan et al., 2020) further empirically verify the scaling law of Transformer architecture in language modeling, regarding the key factors (model size, data size, training cost) and training performance (PPL). Inspired by the scaling law, researchers extend the size of pre-trained language models and further empirically verify the scaling law by training GPT-3 (Brown, 2020). This wave of enthusiasm has led to the creation of GPT-3.5 and GPT-4 (Achiam et al., 2023), ushering in a new era of NLP research and applications.
Based on the scaling laws, the optimal key factors of the model can be determined under given constraints, thus guiding us in model design and resource allocation. Recommendation and advertising systems are mature commercial applications that prioritize ROI, making it highly valuable to explore whether there exists scaling laws for recommendation and advertising models. Due to the lack of a thriving community and open data ecosystem similar to NLP, research on model scaling is relatively scarce in the recommendation and advertisement areas. Early studies primarily gave some qualitative conclusions about model scaling (Shin et al., 2023; Zhang et al., 2024a). Fang et al.((2024)(Fang et al., 2024) first attempt to give a quantitative scaling law of model performance and the amount of training data and model parameters based on public information retrieval benchmarks, and give the practice to solve the optimal amount of data and model parameters under a given total training resource.
However, there is still a lack of literature on scaling law research in real-world online recommendation and advertising systems. We attribute this to two main challenges: 1) For commercial systems, the scaling law should describe the relationship between business revenue and machine costs, rather than the relationship between computing volume, data volume, and offline metrics as seen in traditional research. Achieving this requires training numerous models and conducting extensive online experiments, which are prohibitively expensive in terms of time, manpower, and machine costs; and 2) varying settings across different systems prevent the scaling law from being universally applicable. As a result, to use scaling laws to guide system optimization in a specific scenario, one must first incur substantial costs to obtain the necessary parameters. This makes it impractical for commercial systems to identify and apply scaling laws in many real-world settings.
We aim to identify scaling laws in online advertising systems with low experimental costs and explore their applications in guiding system optimization. An advertising system typically consists of several subsystems, including the advertisement retrieval subsystem and the bidding and ranking subsystem. Each subsystem has its own technology stack, making it challenging to establish a unified paradigm for lightweight identification of scaling laws. In this work, we focus on the online advertisement retrieval subsystem, which selects the top-k ads for the next stage but does not determine the billing of ads. This subsystem is described in detail in Section 3.
To address the issue of high experimental costs, we propose a lightweight paradigm for identifying scaling laws in online advertisement retrieval. First, we introduce a novel offline metric, $R/R^{*}$, which incorporates the revenue information of each ad in the training data and exhibits a high linear correlation with online revenue for retrieval models. Note that we can obtain the specific relationship between $R/R^{*}$ and online revenue for our scenario through the records of daily iterations, theoretically eliminating the need for additional online experiments. Next, we conduct extensive offline experiments to verify the scaling behavior of MLP models in the retrieval scenario, where FLOPs serve as the independent variable and $R/R^{*}$ serves as the dependent variable. Our results show that these scaling behaviors follow a broken neural scaling law (Caballero et al., 2023). Finally, we propose a simple simulation algorithm, developed by our engineering team, to estimate the machine cost. By using the model size parameter as the decision variable and FLOPs and $R/R^{*}$ as intermediate variables, we can associate the model’s estimated machine cost with the estimated online revenue.
We conduct extensive online experiments to verify the effectiveness of our proposed $R/R^{*}$ metric. The results show that $R/R^{*}$ is a more advanced offline surrogate metric for online revenue compared to traditional metrics. When there is no revenue-related information recorded at the ad granularity in the training data, $NDCG$ can serve as a suitable substitute for $R/R^{*}$. Additionally, we demonstrate the applications of the scaling law in ROI-constrained model design and multi-scenario resource allocation. These applications result in a 0.85% and 2.8% improvement in online revenue, respectively, highlighting the significant value of the scaling law in guiding system optimization.
In general, our main contributions are three-fold: 1) We introduce an offline metric, $R/R^{*}$, which exhibits a strong linear correlation with online revenue, bringing significant convenience to the validation of scaling laws and the iterative development of retrieval models. 2) We provide a lightweight paradigm to identify the scaling laws between machine cost and online revenue and verify the existence of scaling laws in industrial advertisement retrieval scenarios. 3) We demonstrate practical applications of the scaling laws in the Kuaishou advertising system. With the scaling law of online revenue and machine cost, we can efficiently perform offline ROI estimations for hundreds of models with different sizes using just a few machines within a few days. The models obtained through this process have led to significant revenue improvement, highlighting the potential of scaling laws in optimizing advertising systems.
2. Related Work
2.1. Neural Scaling Laws
The neural scaling laws, widely recognized in Natural Language Processing and Computer Vision areas, establish predictable relationships between model performance and key factors such as model size, dataset size, computational cost, and post-training error rates. Hestness et al.((2017)(Hestness et al., 2017) introduced a unified scaling law applicable to machine translation, image classification, and speech recognition, noting that different tasks exhibit distinct scaling coefficients. Kaplan et al.((2020)(Kaplan et al., 2020) further elaborated on these laws, defining them through four parameters: model size (number of parameters), dataset size (number of tokens), computational cost (FLOPs), and training loss (perplexity per token). These relationships were empirically validated, including during the training of GPT-3 (Brown, 2020). Subsequently, Hoffmann et al.((2022)(Hoffmann et al., 2022) presented the Chinchilla Scaling Law, which differs somewhat from (Kaplan et al., 2020) because of their different training setups. In addition to estimating the training loss of the model, Isik et al.((2024)(Isik et al., 2024) further verified that there also exist scaling laws between the downstream task metrics and the model parameters. Caballero et al.((2023)(Caballero et al., 2023) found that many scaling behaviors of artificial neural networks follow a smoothly broken power law (BNSL), verified on various NLP and CV tasks, covering a wide range of downstream tasks and model structures.
In the recommendation area, Shin et al.((2023)(Shin et al., 2023) and Zhang et al.((2024a)(Zhang et al., 2024a) conducted studies on whether there exist scaling laws for recommendation models and primarily provided qualitative conclusions. Fang et al.((2024)(Fang et al., 2024) first proposed a quantitative scaling law in the recommendation area, which describes the relationship between the amount of training data, the size of model parameters, and an offline metric for query-document retrieval. Based on this scaling law, the optimal amount of data and model parameter allocation can be solved under a given total training resource. However, obtaining a multivariate scaling law requires a large number of experiments, and the offline metrics might not be a good indicator for online metrics, these make it somewhat difficult to apply in real-world industrial applications. In this paper, we focus on how to obtain the scaling law function between the model’s scalable hyper-parameters (such as the FLOPs) and the online revenue (the primary goal of the advertising system) with only a small amount of experimental cost.
2.2. Models and Evaluation for Online Advertisement Retrieval
Online advertising systems often adopt a cascade ranking framework (Qin et al., 2022; Wang et al., 2011; Gallagher et al., 2019; Chen et al., 2017). The cascade ranking usually includes two types of stages, namely retrieval and ranking. The retrieval stages take a set of terms as input and supply the top-k predicted terms to the next stage, which mainly focuses on order accuracy. The ranking stages in advertising systems should focus on not only the order accuracy but also the calibration accuracy (Huang et al., 2022; Sheng et al., 2023). These lead to some technique differences in retrieval and ranking stages, such as optimization objectives, surrogate losses, design of evaluation metrics, etc. In this work, we focus on the retrieval stages of online advertising systems.
The retrieval stages select the top-k set for the next stage, typically including multiple ranking pathways. In the retrieval stages, which normally refers to the ”Matching” or ”Pre-rank” stage in industrial systems, models often adopt a twin-tower (Covington et al., 2016) or MLP (Hornik et al., 1989) architecture. Some systems may also adopt more complex architectures like transformers (Vaswani et al., 2017). In terms of objective design, industrial systems usually aim to estimate CTR (He et al., 2014) and exposure probability of ads, and use learning-to-rank methods (Wang et al., 2018, 2024) to learn the results of ranking models. For evaluation of retrieval tasks, researchers commonly use OPA, NDCG, and Recall (Wang et al., 2024; Zangerle and Bauer, 2022; Bauer et al., 2024). However, existing metrics often have complex relationships with online revenue, thus we propose $R/R^{*}$, which exhibits a strong linear relationship with online revenue, to reduce the experimental cost of identifying the scaling law for industry scenarios.
3. Formulation of Online Advertisement Retrieval
3.1. Overview of Online Advertising System
Figure 1 illustrates a typical cascade ranking system for online advertisements deployed at Kuaishou Technology. This system is comprised of four main stages: Matching, Pre-ranking, Ranking, and Mix-ranking. The ”Matching” and ”Pre-ranking” stages serve as retrieval mechanisms, focusing on selecting candidate ads without directly impacting their billing. Conversely, the ”Ranking” stage plays a dual role by determining both the selection and billing of ads. It achieves this by predicting the eCPM (Effective Cost Per Mille) of each ad, serving as the standard for billing after its exposure. The ”Mix-ranking” stage integrates the outputs from the advertising and recommendation systems to decide the final set of items presented to the user. For advertising systems, the differences in the goal of the retrieval and ranking stages lead to some technical differences, such as training objectives, system design, evaluation metrics, etc. In this work, we focus on the scaling laws of retrieval stages.
Figure 2 shows the algorithm architecture of the Matching and Pre-ranking stages in the Kuaishou advertising system. They all adopt a multi-pathway ensemble architecture. The Matching stage can be divided into model-based pathways and rule-based pathways. The rule-based pathway can quickly filter out irrelevant ads using predefined criteria, ensuring compliance with regulations and improving ad relevance, while providing flexibility and transparency in the ad selection process. The model-based pathway uses machine learning algorithms to predict and select ads, achieving higher performance ceilings in ad selection. The Pre-ranking stage operates on the set of ads already selected by the Matching stage, and thus only employs the model-based pathways.
Model-based pathways in the Matching and Pre-ranking stages can typically be categorized based on different modeling objectives, such as revenue-oriented, user-interest-oriented, and lifetime value-oriented pathways. In advertising systems, revenue-oriented pathways usually carry the most weight. Since the earlier stages only need to perform set selection, Kuaishou’s advertising system employs Learning-to-Rank (LTR) methods to learn the optimal ranking order to maximize the revenue objective. The LTR pathway is the most prevalent and important in the system, as it is typically the primary focus for algorithm researchers to iterate and improve. Therefore, we use the Pre-ranking LTR pathway as a case study to verify scaling laws in online advertisement retrieval, which similarly applies to Matching models because they share the same multi-pathway architectures.
Figure 1. The cascade ranking architecture at Kuaishou Technology, includes four stages: Matching, Pre-ranking, Ranking, and Mix-ranking. Typically, we regard the ”Matching” and ”Pre-ranking” stages as retrieval stages, which do not decide the billing of ads.
Figure 2. The multi-pathway architecture of Matching and Pre-ranking stages in Kuaishou advertising system. Each pathway has its own weight, which determines the proportion of ads sent to the next stage after the Fusion module.
3.2. Formulation of Retrieval Tasks
Here we give the formulation of retrieval tasks represented by learning-to-rank in the Pre-ranking stage. The organization of training samples and the surrogate loss follows the ARF (Wang et al., 2024). We randomly draw samples from the Pre-ranking, Ranking, and Mix-ranking stages. The training set can be formulated as:
(1) $D_{train}=\{(f_{u_{i}},\{f_{a_{i}^{j}},v_{i}^{j},ecpm_{i}^{j}|1\leq j\leq n\})_{i}\}_{i=1}^{N}$
where $u_{i}$ means the user of the $i$-th impression in the training set, $a^{j}$ means the $j$-th material for ranking in the system. $N$ is the number of impressions of $D_{train}$. The $i$ in $a_{i}^{j}$ means that the sample $a^{j}$ corresponds to impression $i$. The size of the materials for each impression is $n$. $f_{(\cdot)}$ means the feature of ()($\cdot$)( ⋅ ). $v_{i}^{j}$ is the ground truth rank index (the higher value is considered better) of the pair $(u_{i},a^{j})$. The rank index is the relative position of the sampled ads within the system queue, ordered by their original positions. $ecpm_{i}^{j}$ is the eCPM (Effective Cost Per Mille) predicted by the Ranking stage, which can be regarded as the expected revenue for the exposure of $ad_{i}^{j}$. If $ad_{i}^{j}$ does not win in the Pre-ranking and thus does not enter the Ranking stage, the $ecpm_{i}^{j}$ equals 0. The eCPM information in the dataset is only used to construct our offline evaluation metric.
Our learning target is to let the model pick up all the ground-truth ads to the next stage. The surrogate loss of the ARF (Wang et al., 2024) can be formulated as follows:
(2) $L_{total}=L_{Relax}+\frac{1}{2\alpha^{2}}L_{Global}+\log(|\alpha|)$
In Eq 2, $\alpha$ is a trainable scalar, the $L_{Relax}$ and $L_{Global}$ can be formulated as Eq 3 and Eq 4 in the following:
(3) $\begin{split}L_{Relax}&=-\sum_{j}^{n}\{[\sum^{col}\hat{\mathcal{P}}^{\downarrow}_{\textbf{v}_{i}}[:k]]\circ[\log(\frac{1}{m}\sum^{col}\hat{\mathcal{P}}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m])]\}_{j}\end{split}$
(4) $\begin{split}L_{Gobal}&=-\sum_{j}^{n}\sum_{h}^{n}[\hat{\mathcal{P}}^{\downarrow}_{\textbf{v}_{i}}[j,:]\circ\log(\hat{\mathcal{P}}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[j,:])]_{h}\end{split}$
where $\hat{\mathcal{P}}^{\downarrow}_{(\cdot)}$ denote the soft permutation matrices obtained by a differentiable sorting operator (e.g. NeuralSort (Grover et al., 2019)) for sorting vector ()($\cdot$)( ⋅ ) in descending order, $\mathcal{M}^{(i,\cdot)}$ denotes the model predicted score vector of the i-th impression, $\textbf{v}_{i}$ denotes the vector of the ground-truth value for the ads in the i-th impression. $\circ$ refers to the element-wise matrix multiplication operation. $\sum^{col}$ is the column sum for a 2-D matrix. {}$\{\cdot\}_{t}$ and $[\cdot]_{t}$ in Eq 3 and Eq 4 denote the t-th element of the tensor $\cdot$. $m$ and $k$ are hyper-parameters of the ARF.
3.3. Features and Foundation Model
Regarding the features, we adopt both sparse and dense features to describe the information of users and ads in the online advertising system. Sparse features are those whose embeddings are obtained from embedding lookup tables, while dense features are the raw values themselves. The sparse features of the user primarily include the action list of ads and user profile information (e.g., age, gender, and region). The action list mainly consists of action types, frequencies, target ads, and timestamps. The sparse features of the ads primarily include the IDs of the ad and its advertiser. The user’s dense features mainly consist of embeddings produced by other pre-trained models. The dense features of the ads primarily include statistical features information and multimodal features generated by multimodal models.
Regarding the foundation model, we employ the multi-layer perceptrons (MLPs) architecture. The MLP consists of 5 layers, where each hidden layer is composed of batch normalization, linear mapping, and a PReLU (He et al., 2015) activation function in sequence. The output layer is a pure linear layer. Parameters are initialized using the He initialization (He et al., 2015). We apply a log1p transformation as in (Qin et al., 2021) for all statistical dense features. All sparse and dense features are concatenated together and fed as input to the model.
3.4. Background of Online Training and Serving
Our model is trained online using a distributed training framework with synchronous training. We use Adagrad (Duchi et al., 2011) as the optimizer with a learning rate of 0.01 and an epsilon value of 1e-8. The training task consumes data in a streaming manner, approximately processing samples in the order they are generated. The data stream ensures that ads within the same impression are aggregated together and perform a shuffle at the impression level within a small time interval. We use single precision floating point format (FP32) for training and saving the model.
We use a TensorRT-based service architecture to deploy the model for online serving. To accelerate the computation during inference, we adopt half-precision floating point format (FP16) for each layer. Additionally, we decompose the first-layer computation of the MLP into separate user-side and item-side calculations, which are then summed using TensorFlow’s broadcasting mechanism. In this way, the user-side computation needs to be performed only once per impression, while the first-layer item-side computation is cached and periodically refreshed by a dedicated service. We named this strategy ”first-layer optimization”. Since most of the computational cost of an MLP is concentrated in the first layer, the ”first-layer optimization” significantly reduces this cost compared to the original calculation strategy. However, it also introduces a more complex relationship between the model’s FLOPs and the associated machine costs. Due to space limitations, more details about online training and serving are described in appendix A.
4. Scaling Laws of Budget and Revenue
In this section, we verify whether MLPs exhibit scaling laws in the context of online advertisement retrieval, as described in Section 3. We present a lightweight and effective paradigm for identifying the scaling laws of computation budget and revenue. In sub-section 4.1, we introduce $R/R^{*}$, a carefully designed offline metric that is highly linearly correlated with online revenue. In sub-section 4.2, we demonstrate that MLPs exhibit a scaling law between FLOPs and the offline metric $R/R^{*}$. In sub-section 4.3, we propose two computation cost analysis schemes based on model size parameters. Considering the clear relationship between model size parameters and FLOPs, we can approximately determine the expected computing cost and revenue benefits when the model is deployed online with the above work.
4.1. Effective Surrogate Metric for Online Revenue
Traditional offline metrics (e.g. OPA (Wang et al., 2024), Recall, and NDCG) used for online advertising retrieval often ignore the revenue difference of the ground-truth ads and treat them equally. This naturally creates a gap between offline metrics and online revenue. To mitigate this gap, we propose a novel metric named $R/R^{*}$ based on $D_{train}$. The $R/R^{*}_{(i)}$ for impression $i$ in $D_{train}$ can be formulated as Eq 5:
(5) $\begin{split}R/R^{*}_{(i)}=\frac{\sum(\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m;:]*\textbf{ecpm}_{i}^{T})}{\sum(\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}[:m;:]*\textbf{ecpm}_{i}^{T})}\end{split}$
where $\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}$ and $\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}$ denote the hard permutation matrices sorted by $\mathcal{M}^{(i,\cdot)}$ and $\textbf{v}_{i}$ respectively. $\textbf{ecpm}_{i}$ is a row vector that represents the expected revenue of each ad of the $i$-th impression. $R/R^{*}$ is the average of $R/R^{*}_{(i)}$ on $D_{train}$. $m$ is the number of ground-truth ads for exposure.
$R/R^{*}$ aligns better with online revenue mainly because it explicitly considers the commercial value (eCPM) of each ad, directly reflects the goal of maximizing revenue, and reduces the gap between offline evaluation and online performance. If the following two assumptions are met, a linear relationship between $R/R^{*}$ and online revenue is likely to exist:
Assumption 1.
The improvement in $R/R^{*}$ for a single pathway is proportional to the improvement in $R/R^{*}$ of the entire stage ensemble’s output set.
Assumption 2.
The post-stage models generally select only the top-ranked ads from the pre-stage models, and different ads with the same eCPM have the same probability for exposure.
In a mature advertising system, Assumption 2 holds approximately true. Assumption 1 also holds approximately true for pathways with larger weights. Therefore, the $R/R^{*}$ of the LTR pathway and the revenue of the Kuaishou advertising system should be approximately linearly related.
To verify the effectiveness of $R/R^{*}$, we conduct extensive online A/B test experiments. We deploy several different models to empirically study the relationship between online revenue and offline metrics, which differ in feature engineering, surrogate loss, model structure, and model size. Due to the limitations on experimental traffic, we can only run up to 5 A/B test groups simultaneously, which means some experimental groups were launched on different days. The online experiments corresponding to each model use at least 5% of the traffic. We collect each model’s online revenue from the A/B test platform and offline metrics from the training log. Figures 3, 4, 5 and 6 demonstrates the relationship between the online revenue and offline metrics, including our proposed $R/R^{*}$ and traditional retrieval metrics, namely $NDCG$, $Recall$ and $OPA$. To protect commercial secrets, we applied a linear transformation to the original data, and the y-axis represents the relative increase in revenue rather than the absolute value. These operations do not affect the conclusions regarding the linear relationship. Each point in these figures represents data from one day of a model being live. It is evident that:
  • $R/R^{*}$ exhibits the strongest linear correlation with online revenue, with an $R^{2}$ of 0.902, indicating that $R/R^{*}$ is a more advanced offline metric.
  • $NDCG$ exhibits an acceptable linear correlation with online revenue, with an $R^{2}$ of approximately 0.795. Since ads at the same rank in NDCG have the same weight, NDCG is likely a good metric only in systems where ads at the same position have similar eCPM. When training data does not record eCPM information, $NDCG$ would be an alternative to $R/R^{*}$.
  • $Recall$ and $OPA$ exhibit poor linear correlations with online revenue, with $R^{2}$ values of 0.509 and 0.140 respectively, suggesting that they are not good indicators of online revenue.
Figure 3. The relationship between $R/R^{*}$ and the relative increase in online revenue. The $R^{2}$ value is 0.902. The average deviation for the fitted data points is 0.243, and the maximum deviation is 0.483.
Figure 4. The relationship between $NDCG$ and the relative increase in online revenue. The $R^{2}$ value is 0.793. The average deviation for the fitted data points is 0.311, and the maximum deviation is 0.819.
Figure 5. The relationship between $Recall$ and the relative increase in online revenue. The $R^{2}$ value is 0.509. The average deviation for the fitted data points is 0.534, and the maximum deviation is 1.139.
Figure 6. The relationship between $OPA$ and the relative increase in online revenue. The $R^{2}$ value is 0.140. The average deviation for the fitted data points is 0.634, and the maximum deviation is 1.690.
4.2. Scaling Laws of FLOPs and Offline Metric
In this section, we verify whether there is a scaling law for the model between computational effort and online revenue under the setting described in Section 3. Thanks to the strong linear correlation between $R/R^{*}$ and online revenue, we can transform the process of identifying the scaling law from online experiments to offline experiments. We trained 10 models with different MLP sizes offline, each for 10 days to ensure convergence. We then collected their converged $R/R^{*}$ metrics and calculated the FLOPs for each model to verify whether there is a scaling law for MLP models under our setting. We verify whether the collected data follows the Broken Neural Scaling Law (BNSL), which is a statistical law that has been widely validated by Caballero et al.((2023)(Caballero et al., 2023) for many scaling behaviors of artificial neural networks in CV and NLP tasks. BNSL can be formulated as shown in Eq 6:
(6) $\begin{split}R/R^{*}&=a+(b*FLOPs^{-c_{0}})\prod_{i=1}^{t}(1+(\frac{FLOPs}{d_{i}})^{1/f_{i}})^{(-c_{i}*f_{i})}\end{split}$
where $a$,$b$,$c_{0}$,$c_{i}$,$f_{i}$ are parameters and $t$ is hyper-parameter of BNSL. $t$ characterizes the degrees of freedom of BNSL, and in practice, we typically take it as 6. The scaling factor is the FLOPs, and the dependent variable is our proposed offline metric $R/R^{*}$. We use basin-hopping (Wales and Doye, 1997) to fit the BNSL.
Figure 7. Scaling Performance regarding the FLOPs and $R/R^{*}$ of models. The size of the MLP network and the deviation between the fitted curve and the true values (referred to as the ”error”) are annotated in the figure. The $R^{2}$ value of the curve fitting is 0.996. Here [1024, 256, 256, 256, 1] indicates an MLP model with 5 layers, where the output sizes are 1024, 256, 256, 256, and 1, respectively. The total dimensions of input features for all models are 3328.
Figure 7 illustrates the curve obtained after fitting the BNSL, along with the original data points. The x-axis represents the scaling factor (FLOPs), and the y-axis represents the dependent variable ($R/R^{*}$). Each data point corresponds to an MLP model of a different size, with the number of layers and units per layer annotated in the figure. The $R^{2}$ value of the curve fitting is 0.996, indicating a strong correlation between the predicted values and the observed data points. Therefore, it provides strong evidence that there is indeed a Broken Neural Scaling Law for the MLP models under the typical scenario (referred to Section 3) of online advertisement retrieval. With this scaling law, we can predict or understand the scaling behavior of the MLP models.
It is important to note that the FLOPs in Figure 7 refer to the computational cost per user-ad pair (without customized optimization, e.g. the first-layer optimization), and the training data includes about 200 billion samples (namely user-ad pairs) and 20 billion impressions. We set the maximum FLOPs for the model to 143M FLOPs. The FLOPs range is not as wide as in NLP (Kaplan et al., 2020; Hoffmann et al., 2022). This is mainly because 143M FLOPs is about 20 times the FLOPs of our online base model, and a mature commercial system typically cannot easily allocate such a large amount of computing resources in the short term. Additionally, from the scaling laws validated in other scenarios, the marginal returns of increasing computational cost tend to decrease. Therefore, from the perspective of commercial applications of scaling law in advertising systems, which strongly focus on ROI, predicting model performance within the computational range of 0 to 143M FLOPs should be sufficient.
Thanks to the strong linear relationship between $R/R^{*}$ and online revenue, we can establish a prediction from FLOPs to online revenue. To further validate whether the accuracy of the FLOPs-to-revenue estimation meets the needs of model development and iteration, we deploy two models for 7 days, each with 10% of online traffic. The FLOPs for these models are 12.6M and 22.9M, respectively. Based on the scaling law, the estimated online revenue gains are 0.33% and 0.62%, while the actual online revenue increases are 0.38% and 0.69%, respectively. In our scenario, this accuracy is more than sufficient for guiding model development and iteration. Note that the MLP network size is [1024, 512, 512, 512, 1] for both models, while the dimensions of the input features are 5128 and 10128 (calculation details are shown in appendix B.1), respectively. This further indicates the generalization of the FLOPs-to-revenue estimation. We notice that the estimation errors of these two models are smaller than the average error of figure 3. This may be because the models in figure 3 were deployed for only one day, while these two models were deployed for 7 days, which means the observations of figure 3 would have higher variance.
4.3. Mapping Function from Model Settings to Computation Cost and FLOPs
Towards real-world applications, we need the scaling law of revenue and computation cost (namely machine cost), rather than the relationship of revenue and FLOPs. To this end, we should build a robust mapping function from FLOPs to computation cost without online deployment. Unfortunately, it’s intractable to give a certain formulation that maps the FLOPs to computation cost for general scenarios, since the actual computation cost is influenced by a multitude of factors including hardware specifications, software optimization, and so on. These factors can significantly vary across different environments, leading to highly non-linear and unpredictable relationships between FLOPs and actual costs.
To tackle this problem, we employ the parameters of model size to bridge the gap between FLOPs and computation cost. Determining the FLOPs of a model based on its size parameters is trivial. We use the tf.profiler API of TensorFlow to calculate the FLOPs for the forward pass of one user-ad pair. Thus, the main challenge lies in estimating the computation cost based on the model size parameters without online deployment. We propose two paradigms to identify the relationship between machine cost and model size parameters: 1) machine cost estimation based on software and hardware optimization experience, and 2) machine cost estimation based on offline simulation testing.
The first approach relies on professionals with deep expertise in optimizing software and hardware systems to minimize or avoid online deployment operations. They employ empirical methods to establish reliable correlations between model size and actual computation costs. For instance, Fang et al.((2024)(Fang et al., 2024) estimated the machine cost for standard models like Transformers, trained using PyTorch on A100 GPUs, leveraging their expertise. However, in real-world advertising scenarios, customized training and serving frameworks (e.g., the first-layer optimization in section 3.4) and machine configurations introduce more complex nonlinearities between model size parameters and machine cost, increasing the complexity and difficulty of expertise-based cost estimation.
The second method is a more standardized, reproducible, and recommended approach for cost estimation. This method can be formalized as an algorithm, as shown in Algorithm 1. Thanks to the machine cost estimation tools (abbreviated as MCET) developed by our engineering team, which implement Algorithm 1, we can estimate the required machine numbers for tens of models without actually deploying or launching them. Using one machine with a single GPU, we can estimate the required number of machines for a single TensorFlow meta-file within half an hour. Combining this information with the unit price of the specified machine configuration, we can accurately calculate the expected machine cost. It is important to note that for training and serving cost estimation, the algorithm should be executed using the meta files specific to the training and serving process, respectively. For training cost estimation, labels are randomly constructed. This approach not only saves time but also provides a reliable and efficient way to estimate computation costs for model deployment and iteration.
Algorithm 1 Cost Estimation by Offline Simulation Testing
0:  The number $n$ of models (denoted as $n$), and meta files of the models (denoted as $\{meta_{i}|0\leq i<n\}$). One specified machine $\mathcal{MH}$ for online serving. The machine number $req_{0}$ of the online base model $M_{0}$. The desired response time limit $T_{\text{limit}}$.
0:  The estimated machine number of each model for online serving, denoted as $\{req_{i}|1\leq i<n\}$.
1:  Initialize an empty list $results$ to store the results.
2:  Simulate the execution of the base model $M_{0}$ using the specified hardware configuration $\mathcal{MH}$.
3:  Measure the QPS ($QPS_{0}$) of the base model $M_{0}$ under the given response time limit $T_{\text{limit}}$, with the input being randomly initialized, fixed-length tensors.
4:  Calculate the total QPS capacity of the base model setup: $QPS_{\text{total}}=req_{0}\times QPS_{0}$.
5:  for $i=1$ to $n-1$ do
6:     Load the meta file $meta_{i}$ of the $i$-th model.
7:     Simulate the model’s execution using the specified hardware configuration $\mathcal{MH}$.
8:     Measure the QPS ($QPS_{i}$) of the $i$-th model under the given response time limit $T_{\text{limit}}$, with the input being randomly initialized, fixed-length tensors.
9:     Calculate the required number of machines for the $i$-th model: $req_{i}=\frac{QPS_{\text{total}}}{QPS_{i}}$.
10:     Append the result $req_{i}$ to the $results$ list.
11:  end for
12:  return  $results$
5. Applications in Model Designing and Resource Allocation
In this section, we show two real-world applications of the scaling laws identified by the paradigm in section 4 in the advertising system of Kuaishou Technology. To facilitate the description, we first define the following notations:
  • $\mathcal{SP}=[a_{0},a_{1},\ldots,a_{n}]$: A list representing the size parameters of an MLP, where $a_{0}$ denotes the dimension of the MLP input, and $a_{1}$ to $a_{n}$ denote the output dimensions of each layer. Let us denote the TensorFlow meta file generated from the MLP model configured with $\mathcal{SP}$ as $meta_{\mathcal{SP}}$.
  • $F(\mathcal{SP})$: The FLOPs of the network with size parameters $\mathcal{SP}$.
  • $G$: The mapping function from $R/R^{*}$ to the expected online revenue, which is a linear function.
  • $BNSL$: The mapping function from FLOPs to $R/R^{*}$, which follows the form given in Equation 6.
  • $MCET$: The tool (function) used to estimate machine costs, where $MCET(meta_{\mathcal{SP}})$ represents the estimated machine cost for deploying an MLP with size parameters $\mathcal{SP}$.
where $G$, $BNSL$, $F$ and $MCET$ are obtained by referring to section 4.1, section 4.2, and section 4.3, respectively.
5.1. ROI-Constrained Model Designing
For a commercial company, there is usually a standard regarding ROI. Specifically, for an advertising system, we aim for the return on investment (ROI) from the machine to be no less than $\lambda$, where $\lambda$ is typically determined by management based on the company’s operating conditions.
Given a specific MLP model with size parameters $\mathcal{SP}$, the ROI of the model can be estimate by Eq 7:
(7) $\begin{split}ROI&=\frac{G(BNSL(F(\mathcal{SP})))}{MCET(meta_{\mathcal{SP}})}\end{split}$
Thus, given the ROI constraint, we can find the approximate optimal model size by solving the following optimization problem:
(8) $\begin{split}\max_{\mathcal{SP}}&\quad G(BNSL((F(\mathcal{SP}))))\\ \text{s.t. }&\quad\frac{G(BNSL(F(\mathcal{SP})))}{MCET(meta_{\mathcal{SP}})}\geq\lambda\\ \end{split}$
Notice that $\mathcal{SP}$ is a vector where each element represents the number of units in each layer of the MLP. As each element in $\mathcal{SP}$ increases, $F(\mathcal{SP})$, $BNSL(\mathcal{SP})$, and $G(BNSL(\mathcal{SP}))$ also increase, indicating that $F$, $BNSL$, and $G$ are monotonically increasing functions. However, $MCET$ is not a monotonic function and is typically a complex mapping, making the problem non-convex and unsolvable analytically or through binary search.
For the functions in Eq 8, the most time-consuming part is the computation of $MCET(\mathcal{SP})$, which relies on a simulation testing tool. Fortunately, even the computation of $MCET$ does not have a high time and resource cost, allowing us to perform a grid search on this problem. We use 10 T4-GPU machines to run the grid search of about 1000 configurations and find an approximate optimal solution in about two days. Due to space limitations, details of the grid search settings are introduced in appendix B.1.
With the specific ROI constraint of Kuaishou Technology, we solve the approximate optimal size of our foundation MLP model, the $\mathcal{SP}$ of the online base model and the solved model are [3328, 1024, 256, 256, 256, 1] and [16128, 1024, 512, 512, 512, 1], the new model has brought about a 0.85% increase in revenue. Such an improvement is considered significant in our advertising scenario. Without the scaling law, it would be nearly impossible to test the ROI of hundreds of model configurations through online deployment. The time and resource costs of deploying so many models would far outweigh the incremental revenue from the optimized model derived using the scaling law.
5.2. Multi-scenario Resource Allocation
There are often different scenarios that use different models in online advertising systems, such as using different models for different stages and ad placements. Taking different stages as an example, we show the practice of the optimization in multi-scenario resource allocation via scaling law.
Let us assume there are $t$ scenarios, with a total machine cost of $B$. We want to maximize the total revenue through resource allocation across multiple scenarios without increasing the total machine cost, while ensuring that each scenario’s ROI is greater than 1. We can formulate the optimization problem as Eq 9:
(9) $\begin{split}\max_{\{\mathcal{SP}_{i}\mid 0\leq i<t\}}&\sum_{i=0}^{t-1}G(BNSL(F(\mathcal{SP}_{i})))\\ \text{s.t. }&\sum_{i=0}^{t-1}MCET(meta_{\mathcal{SP}_{i}})\leq B\\ \text{s.t. }&\forall i\in\{0,\ldots,t-1\},\frac{G(BNSL(F(\mathcal{SP}_{i})))}{MCET(meta_{\mathcal{SP}_{i}})}\geq 1\\ \end{split}$
This problem can be solved using the same approach as described in sub-section 5.1. The computational complexity of the grid search would be about $t$ times the complexity mentioned in Eq. 8, with the majority of the computation concentrated in the calculation of MCET. In our case, we consider the Matching and Pre-ranking stages as separate scenarios, so $t=2$. Even with $t=2$, we can still use a grid search to solve the problem exhaustively. Due to the initially suboptimal resource allocation, this optimization resulted in a significant 2.8% increase in online revenue. If $t$ is large in practice, making the grid search computationally infeasible, we can employ heuristic information to prune the search space or consider only proportional scaling of the base model. Due to space limitations, more details of multi-scenario resource allocation are shown in appendix B.2.
6. Conclusion
We have systematically studied the scaling laws for online advertisement retrieval and validated the existence of scaling laws in a typical learning-to-rank task with MLP models. To address the challenge of the high cost in time and resources of obtaining scenario-specific scaling law parameters in industrial applications, we propose a lightweight paradigm that transforms most online experiments into low-cost offline experiments through the offline metric R/R* and the machine cost estimation tool. Our work showcases the practical application of the scaling law in optimizing model design under ROI constraints and efficient resource allocation across multiple scenarios in the Kuaishou advertising system, thereby bringing significant commercial value. We also discuss the limitations of our study and outline potential areas for future research in appendix C in the context of scaling laws in online advertisement retrieval.
Acknowledgments
We thank Caiyi Xu, Xiao Liang, Xu He, and Qiang Wang for implementing the machine cost estimation tool and for their daily support of online training and serving issues.
Appendix A More Details about Online Training and Serving
Figure 8 illustrates the pipeline of online training and online serving in the Kuaishou advertising system and the details of ”first-layer optimization” mentioned in section 3.4. The advertising system records logs in real-time for each ad exposure, including the features and labels of the training samples. The downstream training engine reads training data from Kafka storage in real-time and performs online training to optimize the model. Every 20 minutes, the training engine exports the latest model parameters to HDFS for secure storage. Meanwhile, the prediction server periodically polls HDFS to check for new model parameters and updates the local model parameters when a new version is available.
First-layer optimization is an optimization technique designed for MLP models. This technique transforms the computation of the first layer from $FFN(\text{concat}(emb_{user},emb_{ad}))$ (referred to as $FL_{\text{naive}}$) to $FFN(emb_{user})+FFN(emb_{ad})$ (referred to as $FL_{\text{opt}}$). The transformation yields equivalent results but significantly reduces computational overhead. This optimization is applied to the pre-ranking models of the Kuaishou advertising system, where each impression involves evaluating 1500 candidate ads. In this context, for each impression, there is one user embedding ($emb_{user}$) and 1500 ad embeddings ($emb_{ad}$). The operations concat and $+$ are broadcast operations, meaning they can be applied element-wise across arrays of different shapes. Here, the term ”concat” refers to a logical operation with broadcasting properties, which in practice can be implemented using a combination of tf.tile and tf.concat in TensorFlow. By adopting the optimized form $FL_{\text{opt}}$, we can save a substantial amount of computation. To further enhance efficiency, we have developed a scheduled precomputation service called the “Embedding Producer.” This service caches the $FFN(emb_{ad})$ values for all valid ads. It is a CPU-based service that is triggered by model update events and the insertion of new ads. This caching mechanism ensures that the precomputed values are readily available, reducing the computational overhead of the pre-ranking models.
For the training procedure of Pre-ranking models, we have also implemented the first-layer optimization. However, during training, each impression typically involves only about 10 ads, which results in far fewer computational savings compared to the pre-ranking service. For both training and serving, we use T4 GPUs, with each T4 GPU machine equipped with 128 CPU cores. During training, we utilize an asynchronous distributed training framework to improve efficiency and scalability. Both the training and serving frameworks are developed based on TensorFlow. Matching models in our system use the twin-tower structure, inheriting all settings except the first-layer optimization. The inner product of the user and ad tower is directly calculated by brute force for online serving, and no approximate search algorithm such as FAISS is employed.
Figure 8. The pipeline of online training and serving, and details of the first layer optimization for Pre-ranking models.
Appendix B More Details about the Applications of Scaling Law
B.1. ROI-Constrained Model Designing
We set an upper limit on the model FLOPs to 143M and set the step size of the grid search as 16 and 128 for the feature embedding dimensions and the unit number of MLP layers, respectively. This larger step size is chosen because smaller increments have minimal impact on the total FLOPs and result in negligible increases in expected revenue according to the scaling law. The feature embedding dimension ($emb\_dim$) in the grid search starts at 16. The input size of MLP models is determined by multiplying the $emb\_dim$ by the number of sparse features, which is 200 in our experiments, plus an additional 128-dimensional dense feature. For the MLP layers (excluding the output layer, which is fixed at 1), the number of units starts at 128, with the constraint that the number of units in any layer does not exceed the number of units in the previous layer. The size of the first layer outputs does not exceed 1024, because this is the storage limit of the embedding producer. Additionally, we enforce that the number of units in any layer (except the output layer) is at least 1/20 of the number of units in the previous layer as the reason discussed in appendix C.1. We used 10 T4-GPU machines to perform a grid search over approximately 1000 configurations and found an approximate optimal solution in about two days.
B.2. Multi-scenario Resource Allocation
We adopted an MLP model for the Pre-ranking stage and a twin-tower model for the Matching stage. Through offline experiments, we obtained the key functions (including $G$, $BNSL$, etc.) in the scaling laws for both the Matching and Pre-ranking stages using our proposed paradigm. Due to the different model architectures, their machine costs vary in the same size. We used the machine cost estimation tool to additionally measure the machine costs of the twin-tower model at different sizes. Similarly to the estimation of the MLP model, we solved the estimated machine costs of approximately 1000 different sizes of twin-tower models based on the algorithm 1 using 10 T4-GPU machines for about 2 days.
Appendix C Limitations and Future Work
C.1. Optimal Model Designing with Given FLOPs
One remaining question is how to give optimal unit distributions of layers under a given FLOPs. Typically, in advertising models, the size of the bottom layer is the largest, gradually decreasing in the upper layers. We believe that as long as the difference in the number of units between adjacent layers is within a reasonable range, the performance impact should be negligible. To explore this, we conducted preliminary experiments, such as testing a model with the size [3328, 1024, 32, 32, 32, 1]. We found that when the unit allocation is unreasonable, the model performance significantly degrades compared to a reasonable one. Our empirical experience suggests that, except for the output layer, the difference in the number of units between adjacent layers should not exceed 20. Models designed following this principle generally conform to the BNSL. More in-depth conclusions are left for future work and are beyond the scope of this paper.
C.2. New Foundation Models and Joint Laws
While the transformer has become a foundational model in the NLP domain, capable of being applied to almost all tasks and scenarios, the recommendation and advertising fields lack such a foundation model. This absence means that the scaling law parameters obtained in one scenario cannot be readily reused in another. In addition, our scaling laws show that the benefits of increasing FLOPs converge to zero much more rapidly. It’s different from the conclusion in NLP area, where increasing the FLOPs by an order of magnitude typically results in consistent performance improvements. An important reason might be that MLP models in online advertisement retrieval are not as effective foundation models as transformers in NLP. However, it is worth noting that transformers do not universally outperform MLPs in all advertising scenarios, especially when considering the ROI of the machine. Given these challenges, developing more advanced and universal foundation models with better scaling performance (e.g., recent works (Zhai et al., 2024; Zhang et al., 2024b; Guo et al., 2024)) is a crucial direction for future research in online advertisement.
Additionally, obtaining joint laws considering multiple factors such as data, model size, and computational resources is significantly more costly in industrial advertising settings than obtaining single-variable scaling laws. Developing a cost-effective method for deriving joint laws is therefore an important area for future investigation. The high cost of obtaining joint laws is also one of the motivations behind our exploration of unified foundation models.

### v2

@@
 Scaling Laws for Online Advertisement Retrieval
-Abstract.
-The scaling law is a notable property of neural network models that researchers have discovered, which greatly propels the development of large language models.
+Abstract
+The scaling law is a notable property of neural network models and has significantly propelled the development of large language models.
 Scaling laws hold great promise in guiding model design and resource allocation.
-Recent research increasingly shows that scaling laws are not limited to NLP tasks or Transformer architectures; they also apply to domains such as recommendation, where there are scaling laws relating to the amount of training data, model size, and model performance metrics.
-However, there is still a lack of literature on scaling law research in industrial online advertisement retrieval systems.
-This may be because 1) identifying the scaling law for resource cost and online revenue is often expensive in both time and training resources for large-scale industrial applications, and 2) varying settings for different systems prevent the scaling law from being applied across various scenarios.
-To address these issues, we propose a lightweight paradigm to identify the scaling law of online revenue and machine cost for a certain online advertisement retrieval scenario with a low experimental cost.
-Specifically, we focus on a sole factor (FLOPs) and propose an offline metric named $R/R^{*}$ that exhibits a high linear correlation with online revenue for retrieval models.
-We estimate the machine cost offline via a simulation algorithm.
-Thus, we can transform most online experiments into low-cost offline experiments.
-We conduct comprehensive experiments to verify the effectiveness of our proposed metric $R/R^{*}$ and to identify the scaling law in the online advertisement retrieval system of Kuaishou.
-With the scaling law, we demonstrate practical applications for ROI-constrained model designing and multi-scenario resource allocation in the Kuaishou advertising system.
-To the best of our knowledge, this is the first work to study the scaling laws for online advertisement retrieval of real-world large-scale systems, showing great potential for scaling law in advertising system optimization.
-Scaling Law, Online Advertisement Retrieval, Resource Allocation
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning
-1. Introduction
+Recent research increasingly shows that scaling laws are not limited to NLP tasks or Transformer architectures; they also apply to domains such as recommendation.
+However, there is still a lack of literature on scaling law research in online advertisement retrieval systems.
+This may be because 1) identifying the scaling law for resource cost and online revenue is often expensive in both time and training resources for industrial applications, and 2) varying settings for different systems prevent the scaling law from being applied across various scenarios.
+To address these issues, we propose a lightweight paradigm to identify online scaling laws of retrieval models, incorporating a novel offline metric $R/R^{*}$ and an offline simulation algorithm.
+We prove that under mild assumptions, the correlation between $R/R^{*}$ and online revenue asymptotically approaches 1 and empirically validates its effectiveness.
+The simulation algorithm can estimate the machine cost offline.
+Based on the lightweight paradigm, we can identify online scaling laws for retrieval models almost exclusively through offline experiments, and quickly estimate machine costs and revenues for given model configurations.
+We further validate the existence of scaling laws across mainstream model architectures—including Transformer, MLP, and DSSM—in our real-world advertising system.
+With the identified scaling laws, we demonstrate practical applications for ROI-constrained model designing and multi-scenario resource allocation in the online advertising system.
+To the best of our knowledge, this is the first work to study the identification and application of online scaling laws for online advertisement retrieval, showing great potential for scaling laws in advertising system optimization.
+1 Introduction
 The neural scaling laws, describing how neural network performance changes with key factors (e.g.
-model size, dataset size, computational cost), have been discovered in various research areas (Kaplan et al., 2020; Hoffmann et al., 2022; Shin et al., 2023; Isik et al., 2024; Zhang et al., 2024a; Fang et al., 2024).
-Early research (Hestness et al., 2017) shows that the neural network performance is predictable when scaling training data size in various tasks such as neural machine translation and language modeling.
-Kaplan et al.((2020)(Kaplan et al., 2020) further empirically verify the scaling law of Transformer architecture in language modeling, regarding the key factors (model size, data size, training cost) and training performance (PPL).
-Inspired by the scaling law, researchers extend the size of pre-trained language models and further empirically verify the scaling law by training GPT-3 (Brown, 2020).
-This wave of enthusiasm has led to the creation of GPT-3.5 and GPT-4 (Achiam et al., 2023), ushering in a new era of NLP research and applications.
+model size, dataset size, computational cost), have been discovered in various research areas [31, 25, 43, 29, 55, 18].
+Early research [24] shows that the neural network performance is predictable when scaling training data size in various tasks such as neural machine translation and language modeling.
+Kaplan et al.(2020) further empirically verify the scaling law of the Transformer architecture in language modeling, regarding the key factors (model size, data size, training cost) and training performance (PPL).
+Inspired by the scaling law, researchers extend the size of pre-trained language models and further empirically verify the scaling law by training GPT-3 [6].
+This wave of enthusiasm has led to the creation of GPT-3.5 and GPT-4 [3], ushering in a new era of NLP research and applications.
 Based on the scaling laws, the optimal key factors of the model can be determined under given constraints, thus guiding us in model design and resource allocation.
-Recommendation and advertising systems are mature commercial applications that prioritize ROI, making it highly valuable to explore whether there exists scaling laws for recommendation and advertising models.
+Recommendation and advertising systems are mature commercial applications that prioritize ROI, making it highly valuable to explore whether there exist scaling laws for recommendation and advertising models.
 Due to the lack of a thriving community and open data ecosystem similar to NLP, research on model scaling is relatively scarce in the recommendation and advertisement areas.
-Early studies primarily gave some qualitative conclusions about model scaling (Shin et al., 2023; Zhang et al., 2024a).
-Fang et al.((2024)(Fang et al., 2024) first attempt to give a quantitative scaling law of model performance and the amount of training data and model parameters based on public information retrieval benchmarks, and give the practice to solve the optimal amount of data and model parameters under a given total training resource.
-However, there is still a lack of literature on scaling law research in real-world online recommendation and advertising systems.
-We attribute this to two main challenges: 1) For commercial systems, the scaling law should describe the relationship between business revenue and machine costs, rather than the relationship between computing volume, data volume, and offline metrics as seen in traditional research.
-Achieving this requires training numerous models and conducting extensive online experiments, which are prohibitively expensive in terms of time, manpower, and machine costs; and 2) varying settings across different systems prevent the scaling law from being universally applicable.
+Early studies primarily gave some qualitative or offline quantification conclusions about model scaling [43, 55].
+Fang et al.(2024) attempts to give a quantitative scaling law of model performance and the amount of training data and model parameters based on public information retrieval benchmarks, and first gives an offline application practice in the recommendation area, which is to solve the optimal amount of data and model parameters under a given total training resource.
+Crucially, no work has addressed the online identification of scaling laws between business revenue and machine costs in real-world advertising systems.
+We attribute this to two primary challenges: 1) The scaling law must describe the online revenue-cost relationship (not offline metrics), requiring costly online experiments; 2) System-specific configurations prevent universal applicability, making scaling law adoption prohibitively expensive.
 As a result, to use scaling laws to guide system optimization in a specific scenario, one must first incur substantial costs to obtain the necessary parameters.
@@
 We aim to identify scaling laws in online advertising systems with low experimental costs and explore their applications in guiding system optimization.
-An advertising system typically consists of several subsystems, including the advertisement retrieval subsystem and the bidding and ranking subsystem.
-Each subsystem has its own technology stack, making it challenging to establish a unified paradigm for lightweight identification of scaling laws.
-In this work, we focus on the online advertisement retrieval subsystem, which selects the top-k ads for the next stage but does not determine the billing of ads.
+An advertising system typically consists of several subsystems (e.g., advertisement retrieval, bidding, and ranking), each with distinct technical architectures.
+This heterogeneity complicates establishing a unified paradigm for lightweight scaling law identification.
+In this work, we focus on the online advertisement retrieval subsystem—a critical component that selects the top-k ads for downstream stages without directly determining billing.
 This subsystem is described in detail in Section 2.
-To address the issue of high experimental costs, we propose a lightweight paradigm for identifying scaling laws in online advertisement retrieval.
-First, we introduce a novel offline metric, $R/R^{*}$, which incorporates the revenue information of each ad in the training data and exhibits a high linear correlation with online revenue for retrieval models.
-Note that we can obtain the specific relationship between $R/R^{*}$ and online revenue for our scenario through the records of daily iterations, theoretically eliminating the need for additional online experiments.
-Next, we conduct extensive offline experiments to verify the scaling behavior of MLP models in the retrieval scenario, where FLOPs serve as the independent variable and $R/R^{*}$ serves as the dependent variable.
-Our results show that these scaling behaviors follow a broken neural scaling law (Caballero et al., 2023).
-Finally, we propose a simple simulation algorithm, developed by our engineering team, to estimate the machine cost.
-By using the model size parameter as the decision variable and FLOPs and $R/R^{*}$ as intermediate variables, we can associate the model’s estimated machine cost with the estimated online revenue.
-We conduct extensive online experiments to verify the effectiveness of our proposed $R/R^{*}$ metric.
-The results show that $R/R^{*}$ is a more advanced offline surrogate metric for online revenue compared to traditional metrics.
-When there is no revenue-related information recorded at the ad granularity in the training data, $NDCG$ can serve as a suitable substitute for $R/R^{*}$.
-Additionally, we demonstrate the applications of the scaling law in ROI-constrained model design and multi-scenario resource allocation.
-These applications result in a 0.85% and 2.8% improvement in online revenue, respectively, highlighting the significant value of the scaling law in guiding system optimization.
-In general, our main contributions are three-fold: 1) We introduce an offline metric, $R/R^{*}$, which exhibits a strong linear correlation with online revenue, bringing significant convenience to the validation of scaling laws and the iterative development of retrieval models.
-2) We provide a lightweight paradigm to identify the scaling laws between machine cost and online revenue and verify the existence of scaling laws in industrial advertisement retrieval scenarios.
-3) We demonstrate practical applications of the scaling laws in the Kuaishou advertising system.
-With the scaling law of online revenue and machine cost, we can efficiently perform offline ROI estimations for hundreds of models with different sizes using just a few machines within a few days.
-The models obtained through this process have led to significant revenue improvement, highlighting the potential of scaling laws in optimizing advertising systems.
-2. Related Work
-2.1. Neural Scaling Laws
-The neural scaling laws, widely recognized in Natural Language Processing and Computer Vision areas, establish predictable relationships between model performance and key factors such as model size, dataset size, computational cost, and post-training error rates.
-Hestness et al.((2017)(Hestness et al., 2017) introduced a unified scaling law applicable to machine translation, image classification, and speech recognition, noting that different tasks exhibit distinct scaling coefficients.
-Kaplan et al.((2020)(Kaplan et al., 2020) further elaborated on these laws, defining them through four parameters: model size (number of parameters), dataset size (number of tokens), computational cost (FLOPs), and training loss (perplexity per token).
-These relationships were empirically validated, including during the training of GPT-3 (Brown, 2020).
-Subsequently, Hoffmann et al.((2022)(Hoffmann et al., 2022) presented the Chinchilla Scaling Law, which differs somewhat from (Kaplan et al., 2020) because of their different training setups.
-In addition to estimating the training loss of the model, Isik et al.((2024)(Isik et al., 2024) further verified that there also exist scaling laws between the downstream task metrics and the model parameters.
-Caballero et al.((2023)(Caballero et al., 2023) found that many scaling behaviors of artificial neural networks follow a smoothly broken power law (BNSL), verified on various NLP and CV tasks, covering a wide range of downstream tasks and model structures.
-In the recommendation area, Shin et al.((2023)(Shin et al., 2023) and Zhang et al.((2024a)(Zhang et al., 2024a) conducted studies on whether there exist scaling laws for recommendation models and primarily provided qualitative conclusions.
-Fang et al.((2024)(Fang et al., 2024) first proposed a quantitative scaling law in the recommendation area, which describes the relationship between the amount of training data, the size of model parameters, and an offline metric for query-document retrieval.
-Based on this scaling law, the optimal amount of data and model parameter allocation can be solved under a given total training resource.
-However, obtaining a multivariate scaling law requires a large number of experiments, and the offline metrics might not be a good indicator for online metrics, these make it somewhat difficult to apply in real-world industrial applications.
-In this paper, we focus on how to obtain the scaling law function between the model’s scalable hyper-parameters (such as the FLOPs) and the online revenue (the primary goal of the advertising system) with only a small amount of experimental cost.
-2.2. Models and Evaluation for Online Advertisement Retrieval
-Online advertising systems often adopt a cascade ranking framework (Qin et al., 2022; Wang et al., 2011; Gallagher et al., 2019; Chen et al., 2017).
-The cascade ranking usually includes two types of stages, namely retrieval and ranking.
-The retrieval stages take a set of terms as input and supply the top-k predicted terms to the next stage, which mainly focuses on order accuracy.
-The ranking stages in advertising systems should focus on not only the order accuracy but also the calibration accuracy (Huang et al., 2022; Sheng et al., 2023).
-These lead to some technique differences in retrieval and ranking stages, such as optimization objectives, surrogate losses, design of evaluation metrics, etc.
-In this work, we focus on the retrieval stages of online advertising systems.
-The retrieval stages select the top-k set for the next stage, typically including multiple ranking pathways.
-In the retrieval stages, which normally refers to the ”Matching” or ”Pre-rank” stage in industrial systems, models often adopt a twin-tower (Covington et al., 2016) or MLP (Hornik et al., 1989) architecture.
-Some systems may also adopt more complex architectures like transformers (Vaswani et al., 2017).
-In terms of objective design, industrial systems usually aim to estimate CTR (He et al., 2014) and exposure probability of ads, and use learning-to-rank methods (Wang et al., 2018, 2024) to learn the results of ranking models.
-For evaluation of retrieval tasks, researchers commonly use OPA, NDCG, and Recall (Wang et al., 2024; Zangerle and Bauer, 2022; Bauer et al., 2024).
-However, existing metrics often have complex relationships with online revenue, thus we propose $R/R^{*}$, which exhibits a strong linear relationship with online revenue, to reduce the experimental cost of identifying the scaling law for industry scenarios.
-3. Formulation of Online Advertisement Retrieval
-3.1. Overview of Online Advertising System
-Figure 1 illustrates a typical cascade ranking system for online advertisements deployed at Kuaishou Technology.
-This system is comprised of four main stages: Matching, Pre-ranking, Ranking, and Mix-ranking.
-The ”Matching” and ”Pre-ranking” stages serve as retrieval mechanisms, focusing on selecting candidate ads without directly impacting their billing.
-Conversely, the ”Ranking” stage plays a dual role by determining both the selection and billing of ads.
+To address these challenges, we propose a lightweight paradigm for identifying online scaling laws in industrial advertisement retrieval.
+First, we introduce a novel offline metric, $R/R^{*}$, which integrates the predicted revenue of each ad from the training data.
+We prove that, under mild assumptions, the correlation between $R/R^{*}$ and online revenue asymptotically approaches 1.
+This theoretical guarantee, combined with historical A/B test data from daily model iterations, enables us to calibrate the $R/R^{*}$–revenue relationship without requiring additional online experiments.
+Then, we design a simulation algorithm to estimate machine cost from model configurations.
+By treating model size as the decision variable and using FLOPs and $R/R^{*}$ as intermediate proxies, our paradigm enables end-to-end prediction of both machine cost and online revenue, facilitating efficient offline model selection and system optimization.
+We validate the effectiveness of $R/R^{*}$ through extensive online A/B tests.
+Results demonstrate that $R/R^{*}$ serves as a more accurate offline surrogate for online revenue than traditional metrics.
+Then, we conduct extensive offline experiments across mainstream retrieval architectures—Transformer [45], MLP [1], and DSSM [27]—in both the pre-ranking and matching stages.
+Using FLOPs as the independent variable and $R/R^{*}$ as the dependent variable, we observe consistent broken neural scaling laws [9] across all architectures and stages.
+Building on this, we apply the identified scaling laws to two critical use cases: ROI-constrained model design and multi-scenario resource allocation.
+These deployments yield a substantial combined 5.10% improvement in online revenue, demonstrating the practical utility of scaling laws in real-world system optimization.
+Our contributions are threefold: 1) We propose $R/R^{*}$, a novel offline metric that is both theoretically justified (asymptotically correlated with online revenue under mild assumptions) and empirically validated as a superior surrogate for online revenue.
+2) We present a lightweight paradigm for identifying scaling laws between machine cost and online revenue in industrial retrieval systems, enabling offline discovery of broken neural scaling laws across diverse architectures (e.g., Transformer, MLP, DSSM).
+3) We demonstrate real-world applications in ROI-constrained model selection and multi-scenario resource allocation, achieving a substantial 5.10% gain in online revenue.
+Our framework enables rapid offline ROI estimation for hundreds of model configurations within days—accelerating model iteration and system optimization at scale.
+2 Formulation of Online Advertisement Retrieval
+In this section, we clarify the research subject and experimental background of this paper.
+We give a formulation of online advertisement retrieval, including the definition of the retrieval stage, mainstream industrial practice of architecture and learning tasks for retrieval (i.e., multi-pathway architecture and learning-to-rank tasks).
+2.1 Overview of Online Advertising System
+Figure 1 illustrates a typical cascade ranking system for online advertising systems.
+The system is comprised of four main stages: Matching, Pre-ranking, Ranking, and Mix-ranking.
+The “Matching” and “Pre-ranking” stages serve as retrieval mechanisms, focusing on selecting candidate ads without directly impacting their billing.
+Conversely, the “Ranking” stage plays a dual role by determining both the selection and billing of ads.
 It achieves this by predicting the eCPM (Effective Cost Per Mille) of each ad, serving as the standard for billing after its exposure.
-The ”Mix-ranking” stage integrates the outputs from the advertising and recommendation systems to decide the final set of items presented to the user.
+The “Mix-ranking” stage integrates the outputs from the advertising and recommendation systems to decide the final set of items presented to the user.
 For advertising systems, the differences in the goal of the retrieval and ranking stages lead to some technical differences, such as training objectives, system design, evaluation metrics, etc.
 In this work, we focus on the scaling laws of retrieval stages.
-Figure 2 shows the algorithm architecture of the Matching and Pre-ranking stages in the Kuaishou advertising system.
+Figure 2 shows the typical algorithm architecture of the Matching and Pre-ranking stages in advertising systems.
 They all adopt a multi-pathway ensemble architecture.
@@
 In advertising systems, revenue-oriented pathways usually carry the most weight.
-Since the earlier stages only need to perform set selection, Kuaishou’s advertising system employs Learning-to-Rank (LTR) methods to learn the optimal ranking order to maximize the revenue objective.
-The LTR pathway is the most prevalent and important in the system, as it is typically the primary focus for algorithm researchers to iterate and improve.
-Therefore, we use the Pre-ranking LTR pathway as a case study to verify scaling laws in online advertisement retrieval, which similarly applies to Matching models because they share the same multi-pathway architectures.
-Figure 1.
-The cascade ranking architecture at Kuaishou Technology, includes four stages: Matching, Pre-ranking, Ranking, and Mix-ranking.
-Typically, we regard the ”Matching” and ”Pre-ranking” stages as retrieval stages, which do not decide the billing of ads.
-Figure 2.
-The multi-pathway architecture of Matching and Pre-ranking stages in Kuaishou advertising system.
-Each pathway has its own weight, which determines the proportion of ads sent to the next stage after the Fusion module.
-3.2. Formulation of Retrieval Tasks
-Here we give the formulation of retrieval tasks represented by learning-to-rank in the Pre-ranking stage.
-The organization of training samples and the surrogate loss follows the ARF (Wang et al., 2024).
-We randomly draw samples from the Pre-ranking, Ranking, and Mix-ranking stages.
+Since the earlier stages only need to perform set selection, advertising systems often employ Learning-to-rank methods to learn the optimal ranking order to maximize the revenue objective.
+The Learning-to-rank pathway is the most prevalent and important in the system, as it is typically the primary focus for algorithm researchers to iterate and improve [48, 30, 57, 35, 49, 50].
+Therefore, we use the Learning-to-rank pathway as a case study to verify scaling laws in online advertisement retrieval.
+Figure 1: The cascade ranking architecture of our advertising system, includes four stages: Matching, Pre-ranking, Ranking, and Mix-ranking.
+Typically, we regard the “Matching” and “Pre-ranking” stages as retrieval stages, which do not decide the billing of ads.
+Figure 2: A typical multi-pathway architecture of Matching and Pre-ranking stages. Each pathway has its own weight, determining the proportion of ads sent to the next stage after the Fusion module.
+2.2 Formulation of Retrieval Tasks
+Here, we give the formulation of retrieval tasks represented by learning-to-rank.
+Following Zheng et al., Wang et al.(2024, 2025), we randomly draw samples from the Matching, Pre-ranking, Ranking, and Mix-ranking stages.
+We use all the samples for training the model of the Matching stage, and we use the samples from all stages except the Matching stage to train the Pre-ranking models.
 The training set can be formulated as:
-(1) $D_{train}=\{(f_{u_{i}},\{f_{a_{i}^{j}},v_{i}^{j},ecpm_{i}^{j}|1\leq j\leq n\})_{i}\}_{i=1}^{N}$
+$D_{train}=\{(f_{u_{i}},\{f_{a_{i}^{j}},v_{i}^{j},ecpm_{i}^{j}|1\leq j\leq n\})_{i}\}_{i=1}^{N}$ (1)
 where $u_{i}$ means the user of the $i$-th impression in the training set, $a^{j}$ means the $j$-th material for ranking in the system.
@@
 The size of the materials for each impression is $n$.
-$f_{(\cdot)}$ means the feature of ()($\cdot$)( ⋅ ).
-$v_{i}^{j}$ is the ground truth rank index (the higher value is considered better) of the pair $(u_{i},a^{j})$.
+$f_{(\cdot)}$ means the feature of $(\cdot)$.
+$v_{i}^{j}$ is the ground truth rank index (the higher the better) of the pair $(u_{i},a^{j})$.
 The rank index is the relative position of the sampled ads within the system queue, ordered by their original positions.
 $ecpm_{i}^{j}$ is the eCPM (Effective Cost Per Mille) predicted by the Ranking stage, which can be regarded as the expected revenue for the exposure of $ad_{i}^{j}$.
-If $ad_{i}^{j}$ does not win in the Pre-ranking and thus does not enter the Ranking stage, the $ecpm_{i}^{j}$ equals 0.
+If $ad_{i}^{j}$ does not win in the Matching or Pre-ranking and thus does not enter the Ranking stage, the $ecpm_{i}^{j}$ equals 0.
 The eCPM information in the dataset is only used to construct our offline evaluation metric.
-Our learning target is to let the model pick up all the ground-truth ads to the next stage. The surrogate loss of the ARF (Wang et al., 2024) can be formulated as follows:
-(2) $L_{total}=L_{Relax}+\frac{1}{2\alpha^{2}}L_{Global}+\log(|\alpha|)$
-In Eq 2, $\alpha$ is a trainable scalar, the $L_{Relax}$ and $L_{Global}$ can be formulated as Eq 3 and Eq 4 in the following:
-(3) $\begin{split}L_{Relax}&=-\sum_{j}^{n}\{[\sum^{col}\hat{\mathcal{P}}^{\downarrow}_{\textbf{v}_{i}}[:k]]\circ[\log(\frac{1}{m}\sum^{col}\hat{\mathcal{P}}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m])]\}_{j}\end{split}$
-(4) $\begin{split}L_{Gobal}&=-\sum_{j}^{n}\sum_{h}^{n}[\hat{\mathcal{P}}^{\downarrow}_{\textbf{v}_{i}}[j,:]\circ\log(\hat{\mathcal{P}}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[j,:])]_{h}\end{split}$
-where $\hat{\mathcal{P}}^{\downarrow}_{(\cdot)}$ denote the soft permutation matrices obtained by a differentiable sorting operator (e.g.
-NeuralSort (Grover et al., 2019)) for sorting vector ()($\cdot$)( ⋅ ) in descending order, $\mathcal{M}^{(i,\cdot)}$ denotes the model predicted score vector of the i-th impression, $\textbf{v}_{i}$ denotes the vector of the ground-truth value for the ads in the i-th impression.
-$\circ$ refers to the element-wise matrix multiplication operation.
-$\sum^{col}$ is the column sum for a 2-D matrix.
-{}$\{\cdot\}_{t}$ and $[\cdot]_{t}$ in Eq 3 and Eq 4 denote the t-th element of the tensor $\cdot$.
-$m$ and $k$ are hyper-parameters of the ARF.
-3.3. Features and Foundation Model
-Regarding the features, we adopt both sparse and dense features to describe the information of users and ads in the online advertising system.
+The learning objective in Learning-to-Rank is to learn the order of all or part of the training set $D_{train}$.
+This can typically be achieved using different types of methods, such as point-wise [14, 41, 12, 32], pair-wise [10, 58, 7], and list-wise [8, 48, 49, 57] approaches.
+We employ ARF [49] to train the models in our scenario.
+3 Scaling Laws of Budget and Revenue
+In this section, we present a lightweight paradigm for identifying the scaling laws of computation budget (namely, machine cost) and revenue.
+We also verify whether MLP [1], DSSM [27], and Transformer [45] exhibit scaling laws in the context of online advertisement retrieval.
+In sub-section 3.1, we first introduce $R/R^{*}$, a carefully designed offline metric that is highly linearly correlated with online revenue.
+In sub-section 3.2, we then demonstrate that DSSMs, MLPs, and Transformers exhibit a broken neural scaling law [9] between FLOPs and the offline metric $R/R^{*}$.
+In sub-section 3.3, we finally propose a simulation algorithm to estimate machine cost based on model size parameters.
+Leveraging the $R/R^{*}$ metric, the observed scaling laws, and the cost simulation algorithm, we are able to reliably estimate online machine costs and revenue gains from offline model configurations.
+3.1 Effective Surrogate Metric for Online Revenue
+Traditional offline metrics (e.g., OPA [49], Recall, and NDCG) used for online advertising retrieval often ignore the revenue difference of the ground-truth ads and treat them equally.
+This naturally creates a gap between offline metrics and online revenue.
+To mitigate this gap, we propose a novel metric named $R/R^{*}$ based on $D_{train}$.
+The $R/R^{*}_{(i)}$ for impression $i$ in $D_{train}$ can be formulated as Eq 2:
+$\begin{split}R/R_{(i)}^{*}(m)=\frac{\sum(\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m;:]*\textbf{ecpm}_{i}^{T})}{\sum(\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}[:m;:]*\textbf{ecpm}_{i}^{T})}\end{split}$ (2)
+where $\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}$ and $\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}$ denote the hard permutation matrices [20] sorted by $\mathcal{M}^{(i,\cdot)}$ and $\textbf{v}_{i}$ respectively.
+$\mathcal{M}^{(i,\cdot)}$ denotes the prediction vector of model $\mathcal{M}$ for the $i$-th impression of $D_{train}$.
+When the input vector length is $n$, the hard permutation matrix is an $n\times n$ square matrix.
+The notation $[:m;:]$ denotes the operation of taking the first $m$ rows of the matrix.
+$\textbf{ecpm}_{i}$ is a row vector that represents the expected revenue of each ad of the $i$-th impression.
+$R/R^{*}$ is the average of $R/R^{*}_{(i)}$ on $D_{train}$.
+$m$ is a hyper-parameter of $R/R^{*}$.
+$R/R^{*}$ aligns better with online revenue mainly because it explicitly considers the commercial value (eCPM) of each ad, directly reflects the goal of maximizing revenue, and reduces the gap between offline evaluation and online performance.
+If the following assumptions are met, we can prove that the R/R* and online revenue have a linear relationship.
+The proof is in the Appendix B.1.
+Assumption 3.1.
+The training data $D_{train}$ and online data are independent and identically distributed.
+Assumption 3.2.
+The improvement in $R/R^{*}$ for a single pathway is proportional to the improvement in $R/R^{*}$ of the entire stage ensemble’s output set. It can be formulated as:
+$\begin{split}\Delta R/R^{*}_{ensemble}(m)=\alpha\Delta R/R^{*}_{single}(m)\end{split}$ (3)
+where $R/R^{*}_{single}$ are $R/R^{*}_{ensemble}$ the $R/R^{*}$ for single pathway and entire stage, respectively. $\alpha$ is a constant.
+Assumption 3.3.
+Only the top-$m$ ads from the retrieval stages will be selected by the post-stages for exposure; the exposure probability of $ad_{i}^{j}$ is $p_{i}^{j}$.
+We define $\beta_{i}^{j}=\frac{\text{eCPM}_{i}^{j}}{\sum(\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m;:]*\textbf{ecpm}_{i}^{T})}$ as the normalized contribution of the $j$-th ad (on the top-m sets sorted by retrieval stages) of the $i$-th impression.
+We assume that $\beta_{i}^{j}$ and $p_{i}^{j}$ are invariant across different impressions, depending only on the position $j$.
+Assumption 3.4.
+The ranking stage predicts CPMs that are consistent with the true per-impression revenue.
+Moreover, for each impression $i$, the sum of the top-$M$ predicted CPMs (denoted as $\mathcal{P}_{\textbf{v}_{i}}^{\downarrow}[:m;:]*\textbf{ecpm}^{T}$) is statistically independent of the retrieval model’s performance metric $R/R^{*}_{(i)}$.
+In a mature advertising system, we think Assumption ˜3.1, 3.3 and 3.4 holds approximately true.
+Assumption 3.2 also holds approximately true for pathways with dominant weights.
+Therefore, the $R/R^{*}$ of the Learning-to-rank pathway (with dominant weight) and the revenue of online advertising systems should be approximately linearly related.
+To verify the effectiveness of $R/R^{*}$, we also conduct extensive online A/B test experiments.
+We deploy several different models to empirically study the relationship between online revenue and offline metrics, which differ in feature engineering, surrogate loss, model structure, and model size.
+Table 1 shows the experimental results.
+It is evident that $R/R^{*}$ exhibits the strongest linear correlation with online revenue, with an $R^{2}$ of 0.902, indicating that $R/R^{*}$ is a more advanced offline metric.
+Note that $m$ is 2 in the experiment of Table 1.
+Due to space limitations, more experimental details are shown in Appendix Sections˜B.2 and B.3.
+Table 1: Linear correlation analysis between online revenue and four offline ranking metrics across two stages (Pre-ranking and Matching).
+$R^{2}$ measures the goodness of fit (higher is better), while average deviation (Avg.
+Dev) and maximum deviation (Max.
+Dev) reflect the fitting errors (lower is better).
+Best results are bolded.
+Scenario Correlation Measures Offline Metrics
+$R/R^{*}$ NDCG Recall OPA
+Pre-ranking $R^{2}\uparrow$ 0.902 0.793 0.509 0.140
+[…]
+[…]
+Matching $R^{2}\uparrow$ 0.725 0.315 0.504 0.259
+[…]
+[…]
+3.2 Scaling Laws of FLOPs and Offline Metric
+In this section, we investigate whether there exist scaling laws for different model architectures between computational effort and online revenue under the setting described in Section 2.
+Thanks to the strong linear correlation between $R/R^{*}$ and online revenue, we can transform the identification of the scaling laws from costly online A/B tests to efficient offline experiments.
+We train multiple models with different model sizes offline, and then collected their converged $R/R^{*}$ metrics and calculated the FLOPs for each model to examine the existence of scaling laws for mainstream model architectures (i.e., Transformer [45], MLP [1], and DSSM [27]) under two different scenarios (denoted as $Scene_{1}$ and $Scene_{2}$) in our system.
+Specifically, we examine whether the collected data follows the Broken Neural Scaling Law (BNSL), a statistical framework proposed by Caballero et al.(2023) and empirically validated across numerous deep learning tasks in computer vision and natural language processing.
+BNSL is formulated as shown in Eq 4:
+$\begin{split}R/R^{*}&=a+(b*FLOPs^{-c_{0}})\prod_{i=1}^{t}(1+(\frac{FLOPs}{d_{i}})^{1/f_{i}})^{(-c_{i}*f_{i})}\end{split}$
+[…]
+Figure 3: Scaling Performance regarding the FLOPs and $R/R^{*}$ of MLP models under $Scene_{1}$.
+The size of the MLP network and the deviation between the fitted curve and the true values (referred to as the “error”) are annotated in the figure.
+The $R^{2}$ value of the curve fitting is 0.996.
+[1024, 256, 256, 256, 1] indicates an MLP model with 5 layers, where the output sizes are 1024, 256, 256, 256, and 1, respectively.
+The total dimensions of input features for all models are 3328.
+Due to space limitations, we present the results and analysis for MLP models in $Scene_{1}$ as a representative example in the main text.
+Figure 3 illustrates the fitted BNSL curve, along with the original data points.
+The x-axis represents the scaling factor (FLOPs), and the y-axis represents the dependent variable ($R/R^{*}$).
+Each data point corresponds to a model of a different size, with the number of layers and units per layer annotated in the figure.
+The $R^{2}$ value of the curve fitting is 0.996, indicating a strong correlation between the predicted values and the observed data points.
+This provides compelling empirical evidence that a Broken Neural Scaling Law governs the scaling behavior of MLP models in the typical online advertisement retrieval scenario (see Section 2).
+We observe consistent scaling behaviors across other scenarios and architectures.
+Such a scaling law enables us to predict model performance under different computational budgets and better understand efficiency-performance trade-offs.
+Comprehensive experimental setups and results for Transformer, DSSM, and MLP across both $Scene_{1}$ and $Scene_{2}$ are provided in Appendix C, along with an additional discussion on scaling laws in NLP and online advertising systems.
+Thanks to the strong linear relationship between $R/R^{*}$ and online revenue, we can establish a prediction from FLOPs to online revenue.
+To further validate whether the accuracy of the FLOPs-to-revenue estimation meets the needs of model development and iteration, we deploy two MLP models for 7 days, each with 10% of online traffic.
+The FLOPs for these models are $1.26\times 10^{19}$ and $2.29\times 10^{19}$, respectively.
+Based on the scaling law, the estimated online revenue gains are 0.33% and 0.62%, while the actual online revenue increases are 0.38% and 0.69%, respectively.
+In our scenario, this accuracy is more than sufficient for guiding model development and iteration.
+The MLP network has the same size configuration [1024, 512, 512, 512, 1] for both models, while the dimensions of the input features are 5128 and 10128, respectively.
+That is, the scaling law—originally learned by scaling model width under fixed input size—remains accurate even when model scale is increased through dimension expansion of input features, highlighting the robustness and generalization of the FLOPs-to-revenue estimation.
+Notably, the estimation errors of these two models are smaller than the average error of figure 4(a).
+This may be because the models in figure 4(a) were deployed for one day, while these two models were deployed for 7 days, which means the observations of figure 4(a) would have higher variance.
+This observation also suggests the effectiveness and robustness of our proposed offline metric $R/R^{*}$.
+3.3 Mapping Model Settings to Machine Cost
+For real-world model deployment, we need the scaling law of revenue and computation cost (i.e., machine cost), rather than the relationship of revenue and FLOPs.
+Our ultimate goal is to build an end-to-end offline framework that maps model configurations to both revenue and machine cost, enabling ROI-aware model design without online A/B testing.
+The revenue side of this mapping is addressed in Section Sections˜3.1 and 3.2: we establish a robust scaling law from Model $\to$ FLOPs $\to$ $R/R^{*}$ $\to$ Revenue, where FLOPs are analytically computed from model architecture (e.g., using TensorFlow’s tf.profiler).
+However, the cost side cannot be similarly bridged via FLOPs: actual machine cost depends heavily on environment-specific factors—such as hardware, software stack, and system optimizations—making the FLOPs-to-cost relationship highly non-linear and unpredictable across platforms.
+Thus, a direct FLOPs $\to$ Machine Cost mapping is infeasible for general use.
+To overcome this, we introduce model size parameters (e.g., layer dimensions) as a more stable intermediate proxy.
+While FLOPs and model size are closely related, the latter better captures the structural footprint that system-level machine cost estimators rely on.
+Hence, we shift to a Model $\to$ Size Parameters $\to$ Machine Cost pathway, where the key challenge is estimating cost from size without online deployment.
+We discuss two paradigms for this mapping: (1) expert-driven estimation based on system optimization experience, and (2) offline simulation-based estimation.
+The first approach relies on domain experts to empirically correlate model size with machine cost.
+For example, Fang et al.(2024) estimated machine costs for standard Transformer models on A100 GPUs using PyTorch.
+However, in industrial advertising systems, custom training/serving frameworks (e.g., first-layer optimizations in Section D.1) and heterogeneous machine configurations introduce complex, non-standardized dependencies, limiting the generalizability and accuracy of expert-based estimation.
+The second approach—offline simulation—offers a more systematic, reproducible, and scalable solution.
+Given a model architecture and its size parameters, we can generate its meta file offline.
+As detailed in Algorithm 1 (Appendix D.2), our Machine Cost Estimation Tool (MCET) takes such meta files as input and simulates the computational footprint to estimate the required number of machines.
+Using a single GPU, MCET can estimate the machine count for a model within 30 minutes, without actual deployment.
+By combining this with unit machine pricing, we obtain accurate cost projections.
+Notably, for training cost estimation, synthetic labels are used to complete the forward/backward pass simulation.
+This approach enables rapid, reliable machine cost evaluation during model iteration, significantly reducing both time and risk in deployment planning.
+4 Applications in Model Designing and Resource Allocation
+In this section, we demonstrate two real-world applications of the scaling laws established in Section 3: ROI-constrained model design and multi-scenario resource allocation.
+These applications enable cost-aware, data-driven model development without extensive online testing.
+To facilitate exposition, we define the following notations:
+  • $\mathcal{SP}$: The size parameters of a model.
+For MLPs and DSSMs, $\mathcal{SP}=[a_{0},a_{1},\ldots,a_{n}]$ specifies layer dimensions (e.g., input/output sizes).
+For Transformer-based models, $\mathcal{SP}$ includes embedding dimension, number of layers, attention heads, and feed-forward network (FFN) size.
+$meta_{\mathcal{SP}}$ denotes the meta file generated from the model configured with $\mathcal{SP}$.
+  • $F(\mathcal{SP})$: The FLOPs of the model with $\mathcal{SP}$, computed analytically or via profiling tools.
+  • BNSLBNSL: The mapping from FLOPs to $R/R^{*}$, following the form in Equation 4.
+  • $G$: The linear mapping from $R/R^{*}$ to online revenue, calibrated from historical A/B tests.
+  • MCETMCET: Our Machine Cost Estimation Tool, which takes $meta_{\mathcal{SP}}$ as input and simulates system-level resource demands to estimate machine cost, denoted as $MCET(meta_{\mathcal{SP}})$.
+The functions $G$, BNSLBNSL, $F$, and $MCET$ are derived from Sections 3.1, 3.2, and 3.3, respectively.
+4.1 ROI-Constrained Model Designing
+Industrial systems require that any model deployment achieves a minimum return on investment (ROI) threshold $\lambda$, typically set by business stakeholders. We formalize this as:
+$ROI=\frac{G(BNSL(F(\mathcal{SP})))}{MCET(\text{meta}_{\mathcal{SP}})}\geq\lambda$ (5)
+Our goal is to maximize revenue under this constraint:
+$\begin{split}\max_{\mathcal{SP}}&\quad G(BNSL(F(\mathcal{SP})))\\ \text{s.t. }&\quad\frac{G(BNSL(F(\mathcal{SP})))}{MCET(\text{meta}_{\mathcal{SP}})}\geq\lambda\end{split}$ (6)
+While $G\circ BNSL\circ F$ is monotonically increasing in $\mathcal{SP}$, $MCET$ is non-monotonic due to system-level effects (e.g., kernel fusion, memory access patterns), making the problem non-convex.
+Analytical solutions or binary search are infeasible.
+We solve Eq.
+6 via grid search over plausible $\mathcal{SP}$ configurations.
+MCET evaluations are the most time-consuming part, but are still feasible to run offline for grid search.
+Using 10 GPU machines, we evaluate about 1,000 configurations in about two days (details in Appendix D.3).
+Case Study 1: Pre-ranking Model Optimization.
+For the MLP-based pre-ranking model, we identify the optimal $\mathcal{SP}^{*}=[16128,1024,512,512,512,1]$ under our threshold $\lambda^{*}$, improving over the baseline $[3328,1024,256,256,256,1]$.
+This yields a +0.85% online revenue gain.
+Such an improvement is considered significant in our advertising scenario.
+Case Study 2: Matching Pathway Optimization.
+Guided by scaling laws for Transformer models, we introduced a lightweight Transformer-based pathway in the Matching stage, under the constraint of fixed machine cost and $ROI>\lambda^{*}$.
+This change led to a +1.0% increase in overall revenue.
+This approach replaces costly trial-and-error with systematic, offline optimization, enabling rapid iteration on model designs.
+Without the scaling law, it would be nearly impossible to test the ROI of hundreds of model configurations through online deployment.
+The time and resource costs of deploying so many models would far outweigh the incremental revenue from the optimized model derived using the scaling law.
+4.2 Multi-scenario Resource Allocation
+Industrial online advertising systems often deploy different models across scenarios—such as matching, pre-ranking, and ad placements—each with distinct traffic patterns and business objectives.
+To maximize overall revenue under a fixed total machine budget, we propose a scaling-law-driven framework for multi-scenario resource allocation, ensuring $ROI\geq 1$ per scenario.
+Let there be $t$ scenarios with a shared machine budget $B$.
+Our goal is to allocate resources across scenarios to maximize total revenue without exceeding $B$, while maintaining $ROI\geq 1$ in each.
+This leads to the optimization problem in Eq.
+7:
+$\begin{split}\max_{\{\mathcal{SP}_{i}\}_{i=0}^{t-1}}&\;\sum_{i=0}^{t-1}G\big(BNSL(F(\mathcal{SP}_{i}))\big)\\ \text{s.t.
+}&\;\sum_{i=0}^{t-1}MCET(\text{meta}_{\mathcal{SP}_{i}})\leq B\\ &\;\frac{G\big(BNSL(F(\mathcal{SP}_{i}))\big)}{MCET(\text{meta}_{\mathcal{SP}_{i}})}\geq 1,\quad\forall i\in\{0,\ldots,t-1\}\end{split}$ (7)
+where $\mathcal{SP}_{i}$ denotes the model configuration (e.g., layer sizes) in scenario $i$, and $MCET(\cdot)$ estimates machine cost from the model’s meta file.
+This problem can be solved using the same grid search approach as described in sub-section 4.1.
+The computational complexity scales linearly with $t$ compared to the single-scenario case (Eq.
+8), dominated by MCET evaluations.
+Case Study 3: Matching vs.
+Pre-ranking.
+We treat these two stages as separate scenarios ($t=2$).
+Despite the initially suboptimal allocation, our method achieves a +2.8% online revenue gain under the same machine budget, with all scenarios satisfying $ROI\geq 1$.
+Case Study 4: Two Heterogeneous Ad Placements.
+We further validate our framework on two distinct ad placements (i.e., $Scene_{1}$ vs.
+$Scene_{2}$) with different user engagement patterns.
+Applying the same optimization, we observe a +0.45% increase in total revenue, demonstrating the generalizability of our approach across diverse scenario types.
+For large $t$, exhaustive grid search becomes infeasible.
+In such cases, we can employ heuristic search (e.g., Bayesian optimization) or restrict scaling to proportional changes of a base model.
+This framework enables automated, data-driven resource allocation across complex multi-scenario systems, reducing reliance on manual tuning.
+Further details are provided in Appendix D.4.
+5 Conclusion
+We have presented a systematic study of scaling laws in online advertisement retrieval systems, addressing the high cost and inefficiency of online experimentation in industrial settings.
+By introducing an offline metric $R/R^{*}$ and a machine cost estimation method, we propose a lightweight paradigm that transforms costly online trials into efficient offline analyses.
+This paradigm enables accurate modeling of the relationship between machine cost and online revenue—mediated by model configurations—across diverse model architectures.
+We also conduct extensive experiments to validate the existence of scaling laws under different model architectures (i.e., MLP, DSSM, Transformer) and scenarios in an online advertisement retrieval system.
+With our proposed paradigm, we demonstrate practical applications in ROI-constrained model design and multi-scenario resource allocation, achieving a substantial 5.1% improvement in online revenue, validated via A/B testing.
+The framework supports data-driven decision-making and accelerates model iteration in production environments.
+Limitations and future directions are discussed in Appendix E.
+Appendix A Related Work
+A.1 Neural Scaling Laws
+The neural scaling laws, widely recognized in Natural Language Processing and Computer Vision areas, establish predictable relationships between model performance and key factors such as model size, dataset size, computational cost, and post-training error rates.
+Hestness et al.(2017) introduced a unified scaling law applicable to machine translation, image classification, and speech recognition, noting that different tasks exhibit distinct scaling coefficients.
+Kaplan et al.(2020) further elaborated on these laws, defining them through four parameters: model size (number of parameters), dataset size (number of tokens), computational cost (FLOPs), and training loss (perplexity per token).
+These relationships were empirically validated, including during the training of GPT-3 [6].
+Subsequently, Hoffmann et al.(2022) presented the Chinchilla Scaling Law, which differs somewhat from [31] because of their different training setups.
+In addition to estimating the training loss of the model, Isik et al.(2024) further verified that there also exist scaling laws between the downstream task metrics and the model parameters.
+Caballero et al.(2023) found that many scaling behaviors of artificial neural networks follow a smoothly broken power law (BNSL), verified on various NLP and CV tasks, covering a wide range of downstream tasks and model structures.
+In the recommendation area, Shin et al.(2023) and Zhang et al.(2024) studied on whether there exist scaling laws for recommendation models and primarily provided qualitative conclusions.
+Fang et al.(2024) first proposed a quantitative scaling law in the recommendation area, which describes the relationship between the amount of training data, the size of model parameters, and an offline metric for query-document retrieval.
+Based on this scaling law, the optimal amount of data and model parameter allocation can be solved under a given total training resource.
+However, obtaining a multivariate scaling law requires a large number of experiments, and the offline metrics might not be a good indicator for online metrics, these make it somewhat difficult to apply in real-world industrial applications.
+In this paper, we focus on how to obtain the scaling law function between the model’s scalable hyper-parameters (such as the FLOPs) and the online revenue (the primary goal of the advertising system) with only a small amount of experimental cost.
+A.2 Models and Evaluation for Online Advertisement Retrieval
+Online advertising systems often adopt a cascade ranking framework [47, 11, 19, 38, 50].
+The cascade ranking usually includes two types of stages, namely retrieval and ranking.
+The retrieval stages take a set of terms as input and supply the top-k predicted terms to the next stage, which mainly focuses on order accuracy.
+The ranking stages in advertising systems should focus on not only the order accuracy but also the calibration accuracy [28, 42, 56, 15].
+These lead to some technique differences in the retrieval and ranking stages, such as optimization objectives, surrogate losses, and design of evaluation metrics.
+In this work, we focus on the retrieval stages of online advertising systems.
+The retrieval stages select the top-k set for the next stage, typically including multiple ranking pathways.
+In the retrieval stages, which normally refers to the “Matching” or “Pre-rank” stage in industrial systems, the models often adopt a twin-tower [27, 13] or MLP [1, 26, 63, 51, 49, 50] architecture.
+Some systems may also adopt more complex architectures like transformer and its variants [45, 36, 40, 54, 21, 16, 61, 62].
+In terms of objective design, industrial systems usually aim to estimate CTR [23, 59, 60] and exposure probability of ads, and use learning-to-rank methods [7, 48, 30, 49, 50] to learn the results of ranking models.
+For evaluation of retrieval tasks, researchers commonly use OPA, NDCG, and Recall [44, 53, 5, 49, 50].
+However, existing metrics often have complex relationships with online revenue; thus, we propose $R/R^{*}$, which exhibits a strong linear relationship with online revenue, to reduce the experimental cost of identifying the scaling laws for industry scenarios.
+Appendix B More Analysis and Experiment Details of the offline metric $R/R^{*}$
+B.1 Theoretical Analysis on the Relationship between $R/R^{*}$ and Online Revenue
+Here we give the proof that the $R/R^{*}$ metric is linearly correlated with online revenue under the assumptions ˜3.1, 3.2, 3.3 and 3.4 stated in Section 3.1.
+The proof proceeds in two steps: first, we show that the ensemble $R/R^{*}$ is linearly related to revenue on $D_{train}$; second, we show that $R/R^{*}$ of the model of one pathway and the revenue on $D_{train}$ have a linear relationship; finally, we extend this result to online revenue via Assumption 3.1.
+Step 1: The Proof of the Linear Relationship between Ensemble $R/R^{*}$ and Revenue on $D_{train}$.
+Let $p_{i}^{j}$ denote the exposure probability of ad $a_{i}^{j}$ (the likelihood it is selected for display).
+The total revenue can be expressed as:
+$\text{Revenue}=\sum_{i=1}^{N}\sum_{j=1}^{n}p_{i}^{j}\cdot\text{eCPM}_{i}^{j}.$
+By Assumption 3.3, the normalized contribution $\beta_{i}^{j}=\frac{\text{eCPM}_{i}}{\sum(\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m;:]*\textbf{ecpm}_{i}^{T})}$ and $p_{i}^{j}$ depends only on position $j$, denoted as $\beta^{j}$ and $p^{j}$, respectively.
+Let $C_{i}=\sum(\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}[:m;:]*\textbf{ecpm}_{i}^{T})$ (the top-$m$ eCPM sum under ground-truth ranking).
+We rewrite revenue as:
+$\text{Revenue}=\sum_{i=1}^{N}C_{i}\cdot\left(\sum_{j=1}^{n}p^{j}\beta^{j}\cdot\left(\frac{1}{C_{i}}\sum_{j=1}^{m}\mathcal{P}_{\mathcal{M}^{(i,\cdot)}}[:m;:]\cdot\textbf{eCPM}_{i}^{T}\right)\right).$
+Here, $\sum_{j=1}^{m}\mathcal{P}_{\mathcal{M}^{(i,\cdot)}}[:m;:]\cdot\textbf{eCPM}_{i}^{T}$ corresponds to the numerator of $R/R^{*}_{i}(m)$, and $\sum\mathcal{P}_{\textbf{v}_{i}}[:m;:]\cdot\textbf{eCPM}_{i}^{T}$ (that is, $C_{i}$) corresponds to the denominator.
+Thus:
+$\text{Revenue}=\left(\sum_{i=1}^{N}C_{i}\cdot R/R^{*}_{i}(m)\right)\cdot\left(\sum_{j=1}^{n}p^{j}\beta^{j}\right).$
+Under Assumption 3.4, the term $C_{i}$ is independent of $R/R^{*}_{i}$ (since it depends only on the ranking stage’s predictions). Therefore, the revenue becomes:
+$\begin{split}\text{Revenue}&=\left(\sum_{i=1}^{N}C_{i}\right)\cdot\left(\frac{1}{N}\sum_{i=1}^{N}R/R^{*}_{i}(m)\right)\cdot\left(\sum_{j=1}^{n}p^{j}\beta^{j}\right)\\ &=N\cdot\left(\sum_{j=1}^{n}p^{j}\beta^{j}\right)\cdot\mathbb{E}(C)\cdot R/R^{*}(m)\end{split}$ (8)
+This shows that revenue is linearly proportional to the average $R/R^{*}(m)$ over $D_{train}$.
+Note that $R/R^{*}(m)=\frac{1}{N}\sum_{i=1}^{N}R/R^{*}_{i}(m)$, and $\left(\sum_{j=1}^{n}p^{j}\beta^{j}\right)$ can be treat as a constant for different $i$.
+Step 2: Extension to a Single Pathway under the Multi-pathway Architecture. From Step 1, we have established that the ensemble $R/R^{*}_{\text{ensemble}}(m)$ is linearly related to revenue:
+$\text{Revenue}=\gamma\cdot R/R^{*}_{\text{ensemble}}(m),$
+for some constant $\gamma>0$.
+To analyze the relationship between a single pathway and revenue, consider the change in revenue when only pathway $k$ is modified.
+Let $\Delta R/R^{*}_{\text{single},k}(m)$ be the improvement in pathway $k$’s $R/R^{*}$, and assume all other pathways remain unchanged.
+By Assumption 3.2, for a dominant pathway, the corresponding improvement in the ensemble is:
+$\Delta R/R^{*}_{\text{ensemble}}(m)=\alpha\cdot\Delta R/R^{*}_{\text{single},k}(m).$
+Substituting into the revenue equation:
+$\Delta\text{Revenue}=\alpha\cdot\gamma\cdot\Delta R/R^{*}_{\text{single},k}(m).$
+This implies that the change in revenue is directly proportional to the change in $R/R^{*}_{\text{single},k}(m)$, with proportionality constant $\alpha\cdot\gamma$.
+Therefore, for a dominant pathway (where $\omega_{k}\approx 1$), the linear relationship between $R/R^{*}_{\text{single},k}(m)$ and revenue holds.
+This completes the derivation for a single pathway within the multi-pathway architecture.
+Step 3: Extension to Online Revenue.
+By Assumption 3.1, $D_{train}$ and online data are i.i.d.
+Thus, the linear relationship derived on $D_{train}$ directly extends to online revenue.
+Therefore, the linear relationship between $R/R^{*}$ and online revenue holds for individual pathways as well.
+Conclusion: Combining the above results, we conclude that $R/R^{*}$ is linearly correlated with online revenue under the given assumptions.
+This justifies its use as a surrogate metric for offline evaluation.
+B.2 More Experiment Details and the Visualization of the Correlation between Online Revenue and Offline Metrics
+Due to the limitations on experimental traffic, we can only run up to 5 A/B test groups simultaneously, which means some experimental groups were launched on different days.
+The online experiments corresponding to each model use at least 5% of the traffic.
+We collect each model’s online revenue from the A/B test platform and offline metrics from the training log.
+In the main text, we only show some correlation measures (e.g., $R^{2}$) of the linear correlation between online revenue and offline metrics.
+Here we give the visualization results of the original detailed data of online revenue and offline metrics.
+Figure 4 and Figure 5 shows the visualization results for the Pre-ranking and Matching stages of our system, respectively.
+(a) The relationship between $R/R^{*}$ and the relative increase in online revenue. $R^{2}=0.902$, Avg. Dev. = 0.243, Max. Dev. = 0.483.
+(b) The relationship between $NDCG$ and the relative increase in online revenue. $R^{2}=0.793$, Avg. Dev. = 0.311, Max. Dev. = 0.819.
+(c) The relationship between $Recall$ and the relative increase in online revenue. $R^{2}=0.509$, Avg. Dev. = 0.534, Max. Dev. = 1.139.
+(d) The relationship between $OPA$ and the relative increase in online revenue. $R^{2}=0.140$, Avg. Dev. = 0.634, Max. Dev. = 1.690.
+Figure 4: Relationships between evaluation metrics and online revenue in the Pre-ranking stage of our system. Each subplot shows the correlation with $R^{2}$, average deviation, and maximum deviation.
+(a) The relationship between $R/R^{*}$ and the relative increase in online revenue. $R^{2}=0.725$, Avg. Dev. = 0.287, Max. Dev. = 1.082.
+(b) The relationship between $NDCG$ and the relative increase in online revenue. $R^{2}=0.315$, Avg. Dev. = 0.626, Max. Dev. = 1.712.
+(c) The relationship between $Recall$ and the relative increase in online revenue. $R^{2}=0.504$, Avg. Dev. = 0.452, Max. Dev. = 1.100.
+(d) The relationship between $OPA$ and the relative increase in online revenue. $R^{2}=0.259$, Avg. Dev. = 0.558, Max. Dev. = 1.515.
+Figure 5: Relationships between evaluation metrics and online revenue in the Matching stage of our system. Each subplot shows the correlation with $R^{2}$, average deviation, and maximum deviation.
+To protect commercial secrets, we applied a linear transformation to the original data, and the y-axis represents the relative increase in revenue rather than the absolute value.
+These operations do not affect the conclusions regarding the linear relationship.
+Each point in these figures represents data from one day of a model being live.
+Note that all pairwise differences between points are statistically significant ($p<0.05$), which is tested by the online A/B platform.
+We also observe that, in terms of the correlation between offline metrics and online revenue, the pre-ranking stage generally exhibits higher correlation coefficients compared to the matching stage.
+This may be attributed to the fact that, although both stages adopt a multi-pathway architecture, the matching stage involves a larger number of pathways.
+As a result, no single pathway dominates as strongly as in the pre-ranking stage.
+In the multi-pathway architecture, improvements in any individual pathway are more likely to be "covered" or offset by other pathways.
+This observation suggests that, in order to maintain a strong correlation between offline metrics and online performance, it is crucial to preserve dominant pathways with dominant weight within multi-pathway architectures.
+B.3 Hyper-parameter Sensitivity Analysis of $R/R^{*}$
+In offline evaluations, only a limited number of ads are dumped for analysis, with 10 ads sampled per page view (PV, namely impression).
+Among these, 6 slots are reserved for fine-ranked ads, which are the only ones for which eCPM information is available.
+Consequently, we can only assess $R/R^{*}@m$ for $m$ ranging from 1 to 6.
+For the pre-ranking stage, the linear correlation coefficients ($R^{2}$) between these offline metrics and online revenue are reported as follows: $R^{*}(1)=0.891, R/R$$^{*}(2)=0.902, R/R$$^{*}(3)=0.900, R/R$$^{*}(4)=0.892, R/R$$R^{*}(5)=0.887$, and $R/R^{*}(6)=0.830$.
+For the Matching stage, the corresponding linear correlation coefficients ($R^{2}$) are as follows: $R^{*}(1)=0.666, R/R$$^{*}(2)=0.725, R/R$$^{*}(3)=0.699, R/R$$^{*}(4)=0.695, R/R$$R^{*}(5)=0.698$, and $R/R^{*}(6)=0.706$.
+These results indicate that there indeed exists a reasonable $m$ that achieves a high linear correlation with online revenue.
+Specifically, among the evaluated metrics, $R/R^{*}$ demonstrates the highest correlation when $m=2$, with a correlation coefficient of $0.9020$ and $0.725$ for the Pre-ranking and Matching stages, respectively.
+Moreover, it is noteworthy that the $R/R^{*}$ metric incorporates ad revenue information (eCPM) into the calculation of offline metrics, a feature not present in traditional metrics such as NDCG (Normalized Discounted Cumulative Gain) or Recall.
+Traditional metrics primarily focus on ranking accuracy and recall rates without directly accounting for the actual revenue generated by ads.
+Consequently, $R/R^{*}$ exhibits a stronger correlation with revenue across various values of $m$ compared to traditional metrics.
+We believe that the design of the $R/R^{*}$ metric holds significant reference value for the development of offline metrics in cascade ranking systems, especially those aimed at sorting based on a single value such as advertising revenue.
+Appendix C More Details and Discussions of Scaling Laws between FLOPs and $R/R^{*}$
+C.1 Experimental Setup
+C.1.1 Training Setup for MLPs and DSSMs
+Regarding the features: We adopt both sparse and dense features to describe the information of users and ads in the online advertising system.
 Sparse features are those whose embeddings are obtained from embedding lookup tables, while dense features are the raw values themselves.
@@
 The user’s dense features mainly consist of embeddings produced by other pre-trained models.
-The dense features of the ads primarily include statistical features information and multimodal features generated by multimodal models.
-Regarding the foundation model, we employ the multi-layer perceptrons (MLPs) architecture.
-The MLP consists of 5 layers, where each hidden layer is composed of batch normalization, linear mapping, and a PReLU (He et al., 2015) activation function in sequence.
+The dense features of the ads primarily include statistical features information, and multimodal features generated by multimodal models.
+Regarding the MLP models: The MLP consists of 5 layers, where each hidden layer is composed of batch normalization, linear mapping, and a PReLU [22] activation function in sequence.
 The output layer is a pure linear layer.
-Parameters are initialized using the He initialization (He et al., 2015).
-We apply a log1p transformation as in (Qin et al., 2021) for all statistical dense features.
+Parameters are initialized using the He initialization [22].
+We apply a log1p transformation as in [39] for all statistical dense features.
 All sparse and dense features are concatenated together and fed as input to the model.
-3.4. Background of Online Training and Serving
+Regarding the DSSM models: The DSSM consists of a user tower and an item tower.
+Each tower employs an MLP architecture and has 4 layers.
+Each hidden layer is composed of batch normalization, linear mapping, and a PReLU activation function.
+The output layer of the user tower employs only a linear mapping.
+The output layer of the item tower employs a linear mapping and an L2-normalization.
+The initialization strategy and feature processing methods are the same as those described for the MLP model.
 Our model is trained online using a distributed training framework with synchronous training.
-We use Adagrad (Duchi et al., 2011) as the optimizer with a learning rate of 0.01 and an epsilon value of 1e-8.
-The training task consumes data in a streaming manner, approximately processing samples in the order they are generated.
-The data stream ensures that ads within the same impression are aggregated together and perform a shuffle at the impression level within a small time interval.
-We use single precision floating point format (FP32) for training and saving the model.
-We use a TensorRT-based service architecture to deploy the model for online serving.
-To accelerate the computation during inference, we adopt half-precision floating point format (FP16) for each layer.
-Additionally, we decompose the first-layer computation of the MLP into separate user-side and item-side calculations, which are then summed using TensorFlow’s broadcasting mechanism.
-In this way, the user-side computation needs to be performed only once per impression, while the first-layer item-side computation is cached and periodically refreshed by a dedicated service.
-We named this strategy ”first-layer optimization”.
-Since most of the computational cost of an MLP is concentrated in the first layer, the ”first-layer optimization” significantly reduces this cost compared to the original calculation strategy.
-However, it also introduces a more complex relationship between the model’s FLOPs and the associated machine costs.
-Due to space limitations, more details about online training and serving are described in appendix A.
-4. Scaling Laws of Budget and Revenue
-In this section, we verify whether MLPs exhibit scaling laws in the context of online advertisement retrieval, as described in Section 3.
-We present a lightweight and effective paradigm for identifying the scaling laws of computation budget and revenue.
-In sub-section 4.1, we introduce $R/R^{*}$, a carefully designed offline metric that is highly linearly correlated with online revenue.
-In sub-section 4.2, we demonstrate that MLPs exhibit a scaling law between FLOPs and the offline metric $R/R^{*}$.
-In sub-section 4.3, we propose two computation cost analysis schemes based on model size parameters.
-Considering the clear relationship between model size parameters and FLOPs, we can approximately determine the expected computing cost and revenue benefits when the model is deployed online with the above work.
-4.1. Effective Surrogate Metric for Online Revenue
-Traditional offline metrics (e.g.
-OPA (Wang et al., 2024), Recall, and NDCG) used for online advertising retrieval often ignore the revenue difference of the ground-truth ads and treat them equally.
-This naturally creates a gap between offline metrics and online revenue.
-To mitigate this gap, we propose a novel metric named $R/R^{*}$ based on $D_{train}$.
-The $R/R^{*}_{(i)}$ for impression $i$ in $D_{train}$ can be formulated as Eq 5:
-(5) $\begin{split}R/R^{*}_{(i)}=\frac{\sum(\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m;:]*\textbf{ecpm}_{i}^{T})}{\sum(\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}[:m;:]*\textbf{ecpm}_{i}^{T})}\end{split}$
-where $\mathcal{P}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}$ and $\mathcal{P}^{\downarrow}_{\textbf{v}_{i}}$ denote the hard permutation matrices sorted by $\mathcal{M}^{(i,\cdot)}$ and $\textbf{v}_{i}$ respectively.
-$\textbf{ecpm}_{i}$ is a row vector that represents the expected revenue of each ad of the $i$-th impression.
-$R/R^{*}$ is the average of $R/R^{*}_{(i)}$ on $D_{train}$.
-$m$ is the number of ground-truth ads for exposure.
-$R/R^{*}$ aligns better with online revenue mainly because it explicitly considers the commercial value (eCPM) of each ad, directly reflects the goal of maximizing revenue, and reduces the gap between offline evaluation and online performance.
-If the following two assumptions are met, a linear relationship between $R/R^{*}$ and online revenue is likely to exist:
-Assumption 1.
-The improvement in $R/R^{*}$ for a single pathway is proportional to the improvement in $R/R^{*}$ of the entire stage ensemble’s output set.
-Assumption 2.
-The post-stage models generally select only the top-ranked ads from the pre-stage models, and different ads with the same eCPM have the same probability for exposure.
-In a mature advertising system, Assumption 2 holds approximately true.
-Assumption 1 also holds approximately true for pathways with larger weights.
-Therefore, the $R/R^{*}$ of the LTR pathway and the revenue of the Kuaishou advertising system should be approximately linearly related.
-To verify the effectiveness of $R/R^{*}$, we conduct extensive online A/B test experiments.
-We deploy several different models to empirically study the relationship between online revenue and offline metrics, which differ in feature engineering, surrogate loss, model structure, and model size.
-Due to the limitations on experimental traffic, we can only run up to 5 A/B test groups simultaneously, which means some experimental groups were launched on different days.
-The online experiments corresponding to each model use at least 5% of the traffic.
-We collect each model’s online revenue from the A/B test platform and offline metrics from the training log.
-Figures 3, 4, 5 and 6 demonstrates the relationship between the online revenue and offline metrics, including our proposed $R/R^{*}$ and traditional retrieval metrics, namely $NDCG$, $Recall$ and $OPA$.
-To protect commercial secrets, we applied a linear transformation to the original data, and the y-axis represents the relative increase in revenue rather than the absolute value.
-These operations do not affect the conclusions regarding the linear relationship.
-Each point in these figures represents data from one day of a model being live.
-It is evident that:
-  • $R/R^{*}$ exhibits the strongest linear correlation with online revenue, with an $R^{2}$ of 0.902, indicating that $R/R^{*}$ is a more advanced offline metric.
-  • $NDCG$ exhibits an acceptable linear correlation with online revenue, with an $R^{2}$ of approximately 0.795.
-Since ads at the same rank in NDCG have the same weight, NDCG is likely a good metric only in systems where ads at the same position have similar eCPM.
-When training data does not record eCPM information, $NDCG$ would be an alternative to $R/R^{*}$.
-  • $Recall$ and $OPA$ exhibit poor linear correlations with online revenue, with $R^{2}$ values of 0.509 and 0.140 respectively, suggesting that they are not good indicators of online revenue.
-Figure 3.
-The relationship between $R/R^{*}$ and the relative increase in online revenue.
-The $R^{2}$ value is 0.902.
-The average deviation for the fitted data points is 0.243, and the maximum deviation is 0.483.
-Figure 4.
-The relationship between $NDCG$ and the relative increase in online revenue.
-The $R^{2}$ value is 0.793.
-The average deviation for the fitted data points is 0.311, and the maximum deviation is 0.819.
-Figure 5.
-The relationship between $Recall$ and the relative increase in online revenue.
-The $R^{2}$ value is 0.509.
-The average deviation for the fitted data points is 0.534, and the maximum deviation is 1.139.
-Figure 6.
-The relationship between $OPA$ and the relative increase in online revenue.
-The $R^{2}$ value is 0.140.
-The average deviation for the fitted data points is 0.634, and the maximum deviation is 1.690.
-4.2. Scaling Laws of FLOPs and Offline Metric
-In this section, we verify whether there is a scaling law for the model between computational effort and online revenue under the setting described in Section 3.
-Thanks to the strong linear correlation between $R/R^{*}$ and online revenue, we can transform the process of identifying the scaling law from online experiments to offline experiments.
-We trained 10 models with different MLP sizes offline, each for 10 days to ensure convergence.
-We then collected their converged $R/R^{*}$ metrics and calculated the FLOPs for each model to verify whether there is a scaling law for MLP models under our setting.
-We verify whether the collected data follows the Broken Neural Scaling Law (BNSL), which is a statistical law that has been widely validated by Caballero et al.((2023)(Caballero et al., 2023) for many scaling behaviors of artificial neural networks in CV and NLP tasks.
-BNSL can be formulated as shown in Eq 6:
-(6) $\begin{split}R/R^{*}&=a+(b*FLOPs^{-c_{0}})\prod_{i=1}^{t}(1+(\frac{FLOPs}{d_{i}})^{1/f_{i}})^{(-c_{i}*f_{i})}\end{split}$
-where $a$,$b$,$c_{0}$,$c_{i}$,$f_{i}$ are parameters and $t$ is hyper-parameter of BNSL.
-$t$ characterizes the degrees of freedom of BNSL, and in practice, we typically take it as 6.
-The scaling factor is the FLOPs, and the dependent variable is our proposed offline metric $R/R^{*}$.
-We use basin-hopping (Wales and Doye, 1997) to fit the BNSL.
-Figure 7.
-Scaling Performance regarding the FLOPs and $R/R^{*}$ of models.
-The size of the MLP network and the deviation between the fitted curve and the true values (referred to as the ”error”) are annotated in the figure.
-The $R^{2}$ value of the curve fitting is 0.996.
-Here [1024, 256, 256, 256, 1] indicates an MLP model with 5 layers, where the output sizes are 1024, 256, 256, 256, and 1, respectively.
+We use Adagrad [17] as the optimizer with a learning rate of 0.01 and an epsilon value of 1e-8.
+The training dataset comprises approximately 200 billion samples (i.e., user-ad pairs) and 20 billion impressions.
+The largest model, shown in Figure 3, was trained over 10 days using 60 A10 GPUs.
+The training tasks consume a data stream named $D_{\text{stream-train}}$, processing samples in the order they are generated.
+To ensure that all ads within the same impression are processed together for learning-to-rank training, the system groups ads by impression and performs shuffling at the impression level within small time windows.
+C.1.2 Training Setup for Transformers
+To conduct transformer-based experiments, we follow recent practices in generative recommendation [40, 54, 21, 16, 61, 62] and adapt them to our advertising platform.
+We first extract multimodal embeddings using Qwen-VL 2.5 [4].
+To better align these multimodal embeddings with user behavioral signals in recommendation, we fine-tune and transform the embeddings following the QARM framework [34].
+Subsequently, we generate semantic IDs for all items using Res-kmeans.
+We adopt a decoder-only architecture based on the open-source implementation of Qwen3 [52].
+Following Deng et al.
+[16], Zhou et al.
+[61, 62], we treat all impression logs as positive samples and employ full softmax loss for pre-training.
+The pre-trained model is then post-trained using a reward model derived from the online ranking system.
+Specifically, we use the production-grade ad ranking model as the reward model and apply an off-policy reinforcement learning strategy: the post-training data (including rewards) are collected from the live system, rather than generated by the model itself via inference.
+Since the reward model produces list-wise feedback, we employ the ARF [49] loss for post-training.
+On the input side, item features consist solely of semantic IDs.
+For users, the input sequence includes behavioral sequences—such as clicks and conversions—represented by semantic IDs, along with user profile features (e.g., age, gender, context).
+The user profile features are placed at the beginning of the input sequence as a "prompt" to condition the model’s predictions.
+To ensure full model convergence, all Transformer models are trained on live production data for 25 consecutive days, and $R/R^{*}$ is evaluated on the subsequent day’s data.
+For pre-training, we perform incremental batch training using daily-updated datasets, which facilitates efficient data compression similar to the approach in HSTU [54].
+We denote this pre-training dataset as $D_{\text{pre-train}}$.
+For post-training, we initialize from the latest pre-trained checkpoint and perform streaming training using the same data stream (i.e., $D_{\text{stream-train}}$) as described in Appendix C.1.1.
+Due to the significantly larger model size of our largest Transformer compared to MLP and DSSM models, we apply downsampling on $D_{\text{pre-train}}$ and $D_{\text{post-train}}$ to ensure that the daily training data of each scene can be fully processed within 24 hours using 64 flagship GPUs.
+We use the AdamW optimizer with a learning rate of 0.0001, weight decay of 0.01, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\mathrm{e}{-8}$.
+For evaluation, $R/R^{*}$ is computed over $D_{\text{stream-train}}$, where the predicted eCPM for each ad is obtained directly from the reward model’s output score.
+C.2 More Experimental Results of Scaling Laws on Different Model Architectures
+In addition to the key results presented in the main text, we provide extended experimental validation of scaling laws across multiple model architectures and system stages.
+We examine MLP and DSSM models across two business scenarios, confirming consistent scaling behavior between FLOPs and $R/R^{*}$.
+We further validate the existence of such scaling laws for Transformer-based architectures in the same industrial advertisement retrieval settings.
+Here, we present additional results for the pre-ranking stage in $Scene_{2}$ using the MLP architecture, as shown in Figure 6.
+We also include results for the matching stage in $Scene_{1}$ and $Scene_{2}$ using the DSSM model, corresponding to Figure 7 and Figure 8, respectively.
+For Transformer-based models, we show the scaling curves on the matching stage in $Scene_{1}$ and $Scene_{2}$ in Figure 9 and Figure 10, respectively.
+These results further support the existence of a broken neural scaling law in real-world industrial retrieval systems across diverse architectures.
+Importantly, this does not imply that the BNSL parameters are identical across architectures; rather, it means that within each architectural family, performance scales predictably with FLOPs in the observed FLOPs range.
+Figure 6: Scaling Performance regarding the FLOPs and $R/R^{*}$ of MLP models under $Scene_{2}$.
+The size of the MLP network and the deviation between the fitted curve and the true values (referred to as the “error”) are annotated in the figure.
+The $R^{2}$ value of the curve fitting is 0.992.
+[1024, 512, 512, 1] indicates an MLP model with 4 layers, where the output sizes are 1024, 512, 512, and 1, respectively.
 The total dimensions of input features for all models are 3328.
-Figure 7 illustrates the curve obtained after fitting the BNSL, along with the original data points.
-The x-axis represents the scaling factor (FLOPs), and the y-axis represents the dependent variable ($R/R^{*}$).
-Each data point corresponds to an MLP model of a different size, with the number of layers and units per layer annotated in the figure.
-The $R^{2}$ value of the curve fitting is 0.996, indicating a strong correlation between the predicted values and the observed data points.
-Therefore, it provides strong evidence that there is indeed a Broken Neural Scaling Law for the MLP models under the typical scenario (referred to Section 3) of online advertisement retrieval.
-With this scaling law, we can predict or understand the scaling behavior of the MLP models.
-It is important to note that the FLOPs in Figure 7 refer to the computational cost per user-ad pair (without customized optimization, e.g.
-the first-layer optimization), and the training data includes about 200 billion samples (namely user-ad pairs) and 20 billion impressions.
-We set the maximum FLOPs for the model to 143M FLOPs.
-The FLOPs range is not as wide as in NLP (Kaplan et al., 2020; Hoffmann et al., 2022).
-This is mainly because 143M FLOPs is about 20 times the FLOPs of our online base model, and a mature commercial system typically cannot easily allocate such a large amount of computing resources in the short term.
-Additionally, from the scaling laws validated in other scenarios, the marginal returns of increasing computational cost tend to decrease.
-Therefore, from the perspective of commercial applications of scaling law in advertising systems, which strongly focus on ROI, predicting model performance within the computational range of 0 to 143M FLOPs should be sufficient.
-Thanks to the strong linear relationship between $R/R^{*}$ and online revenue, we can establish a prediction from FLOPs to online revenue.
-To further validate whether the accuracy of the FLOPs-to-revenue estimation meets the needs of model development and iteration, we deploy two models for 7 days, each with 10% of online traffic.
-The FLOPs for these models are 12.6M and 22.9M, respectively.
-Based on the scaling law, the estimated online revenue gains are 0.33% and 0.62%, while the actual online revenue increases are 0.38% and 0.69%, respectively.
-In our scenario, this accuracy is more than sufficient for guiding model development and iteration.
-Note that the MLP network size is [1024, 512, 512, 512, 1] for both models, while the dimensions of the input features are 5128 and 10128 (calculation details are shown in appendix B.1), respectively.
-This further indicates the generalization of the FLOPs-to-revenue estimation.
-We notice that the estimation errors of these two models are smaller than the average error of figure 3.
-This may be because the models in figure 3 were deployed for only one day, while these two models were deployed for 7 days, which means the observations of figure 3 would have higher variance.
-4.3. Mapping Function from Model Settings to Computation Cost and FLOPs
-Towards real-world applications, we need the scaling law of revenue and computation cost (namely machine cost), rather than the relationship of revenue and FLOPs.
-To this end, we should build a robust mapping function from FLOPs to computation cost without online deployment.
-Unfortunately, it’s intractable to give a certain formulation that maps the FLOPs to computation cost for general scenarios, since the actual computation cost is influenced by a multitude of factors including hardware specifications, software optimization, and so on.
-These factors can significantly vary across different environments, leading to highly non-linear and unpredictable relationships between FLOPs and actual costs.
-To tackle this problem, we employ the parameters of model size to bridge the gap between FLOPs and computation cost.
-Determining the FLOPs of a model based on its size parameters is trivial.
-We use the tf.profiler API of TensorFlow to calculate the FLOPs for the forward pass of one user-ad pair.
-Thus, the main challenge lies in estimating the computation cost based on the model size parameters without online deployment.
-We propose two paradigms to identify the relationship between machine cost and model size parameters: 1) machine cost estimation based on software and hardware optimization experience, and 2) machine cost estimation based on offline simulation testing.
-The first approach relies on professionals with deep expertise in optimizing software and hardware systems to minimize or avoid online deployment operations.
-They employ empirical methods to establish reliable correlations between model size and actual computation costs.
-For instance, Fang et al.((2024)(Fang et al., 2024) estimated the machine cost for standard models like Transformers, trained using PyTorch on A100 GPUs, leveraging their expertise.
-However, in real-world advertising scenarios, customized training and serving frameworks (e.g., the first-layer optimization in section 3.4) and machine configurations introduce more complex nonlinearities between model size parameters and machine cost, increasing the complexity and difficulty of expertise-based cost estimation.
-The second method is a more standardized, reproducible, and recommended approach for cost estimation.
-This method can be formalized as an algorithm, as shown in Algorithm 1.
-Thanks to the machine cost estimation tools (abbreviated as MCET) developed by our engineering team, which implement Algorithm 1, we can estimate the required machine numbers for tens of models without actually deploying or launching them.
-Using one machine with a single GPU, we can estimate the required number of machines for a single TensorFlow meta-file within half an hour.
-Combining this information with the unit price of the specified machine configuration, we can accurately calculate the expected machine cost.
-It is important to note that for training and serving cost estimation, the algorithm should be executed using the meta files specific to the training and serving process, respectively.
-For training cost estimation, labels are randomly constructed.
-This approach not only saves time but also provides a reliable and efficient way to estimate computation costs for model deployment and iteration.
-Algorithm 1 Cost Estimation by Offline Simulation Testing
-0:  The number $n$ of models (denoted as $n$), and meta files of the models (denoted as $\{meta_{i}|0\leq i<n\}$).
-One specified machine $\mathcal{MH}$ for online serving.
-The machine number $req_{0}$ of the online base model $M_{0}$.
-The desired response time limit $T_{\text{limit}}$.
-0:  The estimated machine number of each model for online serving, denoted as $\{req_{i}|1\leq i<n\}$.
-1:  Initialize an empty list $results$ to store the results.
-2:  Simulate the execution of the base model $M_{0}$ using the specified hardware configuration $\mathcal{MH}$.
-3:  Measure the QPS ($QPS_{0}$) of the base model $M_{0}$ under the given response time limit $T_{\text{limit}}$, with the input being randomly initialized, fixed-length tensors.
-4:  Calculate the total QPS capacity of the base model setup: $QPS_{\text{total}}=req_{0}\times QPS_{0}$.
-5:  for $i=1$ to $n-1$ do
-6:     Load the meta file $meta_{i}$ of the $i$-th model.
-7:     Simulate the model’s execution using the specified hardware configuration $\mathcal{MH}$.
-8:     Measure the QPS ($QPS_{i}$) of the $i$-th model under the given response time limit $T_{\text{limit}}$, with the input being randomly initialized, fixed-length tensors.
-9:     Calculate the required number of machines for the $i$-th model: $req_{i}=\frac{QPS_{\text{total}}}{QPS_{i}}$.
-10:     Append the result $req_{i}$ to the $results$ list.
-11:  end for
-12:  return  $results$
-5. Applications in Model Designing and Resource Allocation
-In this section, we show two real-world applications of the scaling laws identified by the paradigm in section 4 in the advertising system of Kuaishou Technology.
-To facilitate the description, we first define the following notations:
-  • $\mathcal{SP}=[a_{0},a_{1},\ldots,a_{n}]$: A list representing the size parameters of an MLP, where $a_{0}$ denotes the dimension of the MLP input, and $a_{1}$ to $a_{n}$ denote the output dimensions of each layer.
-Let us denote the TensorFlow meta file generated from the MLP model configured with $\mathcal{SP}$ as $meta_{\mathcal{SP}}$.
-  • $F(\mathcal{SP})$: The FLOPs of the network with size parameters $\mathcal{SP}$.
-  • $G$: The mapping function from $R/R^{*}$ to the expected online revenue, which is a linear function.
-  • $BNSL$: The mapping function from FLOPs to $R/R^{*}$, which follows the form given in Equation 6.
-  • $MCET$: The tool (function) used to estimate machine costs, where $MCET(meta_{\mathcal{SP}})$ represents the estimated machine cost for deploying an MLP with size parameters $\mathcal{SP}$.
-where $G$, $BNSL$, $F$ and $MCET$ are obtained by referring to section 4.1, section 4.2, and section 4.3, respectively.
-5.1. ROI-Constrained Model Designing
-For a commercial company, there is usually a standard regarding ROI.
-Specifically, for an advertising system, we aim for the return on investment (ROI) from the machine to be no less than $\lambda$, where $\lambda$ is typically determined by management based on the company’s operating conditions.
-Given a specific MLP model with size parameters $\mathcal{SP}$, the ROI of the model can be estimate by Eq 7:
-(7) $\begin{split}ROI&=\frac{G(BNSL(F(\mathcal{SP})))}{MCET(meta_{\mathcal{SP}})}\end{split}$
-Thus, given the ROI constraint, we can find the approximate optimal model size by solving the following optimization problem:
-(8) $\begin{split}\max_{\mathcal{SP}}&\quad G(BNSL((F(\mathcal{SP}))))\\ \text{s.t. }&\quad\frac{G(BNSL(F(\mathcal{SP})))}{MCET(meta_{\mathcal{SP}})}\geq\lambda\\ \end{split}$
-Notice that $\mathcal{SP}$ is a vector where each element represents the number of units in each layer of the MLP.
-As each element in $\mathcal{SP}$ increases, $F(\mathcal{SP})$, $BNSL(\mathcal{SP})$, and $G(BNSL(\mathcal{SP}))$ also increase, indicating that $F$, $BNSL$, and $G$ are monotonically increasing functions.
-However, $MCET$ is not a monotonic function and is typically a complex mapping, making the problem non-convex and unsolvable analytically or through binary search.
-For the functions in Eq 8, the most time-consuming part is the computation of $MCET(\mathcal{SP})$, which relies on a simulation testing tool.
-Fortunately, even the computation of $MCET$ does not have a high time and resource cost, allowing us to perform a grid search on this problem.
-We use 10 T4-GPU machines to run the grid search of about 1000 configurations and find an approximate optimal solution in about two days.
-Due to space limitations, details of the grid search settings are introduced in appendix B.1.
-With the specific ROI constraint of Kuaishou Technology, we solve the approximate optimal size of our foundation MLP model, the $\mathcal{SP}$ of the online base model and the solved model are [3328, 1024, 256, 256, 256, 1] and [16128, 1024, 512, 512, 512, 1], the new model has brought about a 0.85% increase in revenue.
-Such an improvement is considered significant in our advertising scenario.
-Without the scaling law, it would be nearly impossible to test the ROI of hundreds of model configurations through online deployment.
-The time and resource costs of deploying so many models would far outweigh the incremental revenue from the optimized model derived using the scaling law.
-5.2. Multi-scenario Resource Allocation
-There are often different scenarios that use different models in online advertising systems, such as using different models for different stages and ad placements.
-Taking different stages as an example, we show the practice of the optimization in multi-scenario resource allocation via scaling law.
-Let us assume there are $t$ scenarios, with a total machine cost of $B$.
-We want to maximize the total revenue through resource allocation across multiple scenarios without increasing the total machine cost, while ensuring that each scenario’s ROI is greater than 1.
-We can formulate the optimization problem as Eq 9:
-(9) $\begin{split}\max_{\{\mathcal{SP}_{i}\mid 0\leq i<t\}}&\sum_{i=0}^{t-1}G(BNSL(F(\mathcal{SP}_{i})))\\ \text{s.t.
-}&\sum_{i=0}^{t-1}MCET(meta_{\mathcal{SP}_{i}})\leq B\\ \text{s.t.
-}&\forall i\in\{0,\ldots,t-1\},\frac{G(BNSL(F(\mathcal{SP}_{i})))}{MCET(meta_{\mathcal{SP}_{i}})}\geq 1\\ \end{split}$
-This problem can be solved using the same approach as described in sub-section 5.1.
-The computational complexity of the grid search would be about $t$ times the complexity mentioned in Eq.
-8, with the majority of the computation concentrated in the calculation of MCET.
-In our case, we consider the Matching and Pre-ranking stages as separate scenarios, so $t=2$.
-Even with $t=2$, we can still use a grid search to solve the problem exhaustively.
-Due to the initially suboptimal resource allocation, this optimization resulted in a significant 2.8% increase in online revenue.
-If $t$ is large in practice, making the grid search computationally infeasible, we can employ heuristic information to prune the search space or consider only proportional scaling of the base model.
-Due to space limitations, more details of multi-scenario resource allocation are shown in appendix B.2.
-6. Conclusion
-We have systematically studied the scaling laws for online advertisement retrieval and validated the existence of scaling laws in a typical learning-to-rank task with MLP models.
-To address the challenge of the high cost in time and resources of obtaining scenario-specific scaling law parameters in industrial applications, we propose a lightweight paradigm that transforms most online experiments into low-cost offline experiments through the offline metric R/R* and the machine cost estimation tool.
-Our work showcases the practical application of the scaling law in optimizing model design under ROI constraints and efficient resource allocation across multiple scenarios in the Kuaishou advertising system, thereby bringing significant commercial value.
-We also discuss the limitations of our study and outline potential areas for future research in appendix C in the context of scaling laws in online advertisement retrieval.
-Acknowledgments
-We thank Caiyi Xu, Xiao Liang, Xu He, and Qiang Wang for implementing the machine cost estimation tool and for their daily support of online training and serving issues.
-Appendix A More Details about Online Training and Serving
-Figure 8 illustrates the pipeline of online training and online serving in the Kuaishou advertising system and the details of ”first-layer optimization” mentioned in section 3.4.
+Figure 7: Scaling Performance regarding the FLOPs and $R/R^{*}$ of DSSM models under $Scene_{1}$.
+The size of the tower in DSSM and the deviation between the fitted curve and the true values (referred to as the “error”) are annotated in the figure.
+The $R^{2}$ value of the curve fitting is 0.998.
+[1024, 512, 512, 64] indicates each tower of DSSM with 4 layers, where the output sizes are 1024, 512, 512, and 64, respectively.
+The total dimensions of input features for the user and item towers are 1760 and 1568, respectively.
+Figure 8: Scaling Performance regarding the FLOPs and $R/R^{*}$ of DSSM models under $Scene_{2}$.
+The size of the tower in DSSM and the deviation between the fitted curve and the true values (referred to as the “error”) are annotated in the figure.
+The $R^{2}$ value of the curve fitting is 0.999.
+[1024, 512, 512, 64] indicates each tower of DSSM with 4 layers, where the output sizes are 1024, 512, 512, and 64, respectively.
+The total dimensions of input features for the user and item towers are 1760 and 1568, respectively.
+Figure 9: Scaling Performance regarding the FLOPs and $R/R^{*}$ of Transformer models under $Scene_{1}$.
+The size of Transformer models and the deviation between the fitted curve and the true values (referred to as the “error”) are annotated in the figure.
+The $R^{2}$ value of the curve fitting is 0.984.
+Figure 10: Scaling Performance regarding the FLOPs and $R/R^{*}$ of Transformer models under $Scene_{2}$.
+The size of Transformer models and the deviation between the fitted curve and the true values (referred to as the “error”) are annotated in the figure.
+The $R^{2}$ value of the curve fitting is 0.992.
+C.3 Discussion: Comparison with Scaling Laws in NLP
+The scaling behaviors observed in our industrial advertisement retrieval system share formal similarities with those in NLP: both exhibit power-law relationships between model size (or FLOPs) and performance within certain computational ranges, enabling predictable performance trends.
+However, the practical application, scope, and implications of scaling laws differ significantly due to the distinct operational constraints of advertising systems.
+A key difference lies in the direction of extrapolation.
+In NLP, scaling laws are typically used to extrapolate upward—predicting large-scale performance from small-scale experiments.
+In contrast, in advertising retrieval, it is often feasible to run large models on small traffic slices and then infer their behavior at smaller, deployable scales.
+This "downward inference" is enabled by the ability to safely test high-cost models offline or on limited traffic, making scaling laws a practical tool for ROI-driven model selection.
+Another critical distinction is the range of computation explored.
+As exemplified by our experiments on the MLP architecture, the maximum FLOPs tested ($1.43\times 10^{20}$) is approximately 20 times that of our online base model.
+While this range is far narrower than those in large-scale NLP studies [31, 25], it is sufficient for industrial deployment.
+This limitation arises not from methodological constraints, but from the economic reality of ROI optimization: beyond a certain point, the diminishing marginal returns implied by power-law scaling make further increases in model size unjustifiable.
+Even if performance continues to improve, the cost-benefit ratio becomes unfavorable.
+Therefore, in the advertising domain, the value of scaling laws lies not in enabling ever-larger models, but in guiding efficient resource allocation and model design within practical budgets.
+We argue that future breakthroughs will likely stem not from scaling within existing paradigms, but from architectural or foundational innovations—what we term paradigm innovations—that shift the scaling curve itself.
+This perspective is further discussed in Appendix E.
+Appendix D More Details about the Applications of Scaling Law
+D.1 Details about Infra Settings for Online Training and Serving
+Figure 11 illustrates the pipeline of online training and online serving in our advertising system and the details of “first-layer optimization” mentioned in section 3.3.
 The advertising system records logs in real-time for each ad exposure, including the features and labels of the training samples.
@@
 The transformation yields equivalent results but significantly reduces computational overhead.
-This optimization is applied to the pre-ranking models of the Kuaishou advertising system, where each impression involves evaluating 1500 candidate ads.
+This optimization is applied to the pre-ranking models of our advertising system, where each impression involves evaluating 1500 candidate ads.
 In this context, for each impression, there is one user embedding ($emb_{user}$) and 1500 ad embeddings ($emb_{ad}$).
-The operations concat and $+$ are broadcast operations, meaning they can be applied element-wise across arrays of different shapes.
-Here, the term ”concat” refers to a logical operation with broadcasting properties, which in practice can be implemented using a combination of tf.tile and tf.concat in TensorFlow.
+The operations concat and ++ are broadcast operations, meaning they can be applied element-wise across arrays of different shapes.
+Here, the term “concat” refers to a logical operation with broadcasting properties, which in practice can be implemented using a combination of tf.tile and tf.concat in TensorFlow.
 By adopting the optimized form $FL_{\text{opt}}$, we can save a substantial amount of computation.
@@
 This caching mechanism ensures that the precomputed values are readily available, reducing the computational overhead of the pre-ranking models.
+However, it also introduces a more complex relationship between the model’s FLOPs and the associated machine costs.
+We use a TensorRT-based service architecture to deploy the model for online serving.
+To accelerate the computation during inference, we adopt half-precision floating point format (FP16) for each layer.
 For the training procedure of Pre-ranking models, we have also implemented the first-layer optimization.
-However, during training, each impression typically involves only about 10 ads, which results in far fewer computational savings compared to the pre-ranking service.
-For both training and serving, we use T4 GPUs, with each T4 GPU machine equipped with 128 CPU cores.
-During training, we utilize an asynchronous distributed training framework to improve efficiency and scalability.
-Both the training and serving frameworks are developed based on TensorFlow.
-Matching models in our system use the twin-tower structure, inheriting all settings except the first-layer optimization.
+We use single precision floating point format (FP32) for training and saving the MLP and DSSM models.
+We use FP16 for training Transformer models.
+For training and serving DSSM and Transformer models in Matching stages, we incorporate several standard engineering optimizations to improve efficiency.
+These include model quantization, user tower computation compression, and caching of precomputed item tower embeddings.
+These techniques significantly reduce both the latency and the overall computational overhead during online training and serving.
 The inner product of the user and ad tower is directly calculated by brute force for online serving, and no approximate search algorithm such as FAISS is employed.
-Figure 8. The pipeline of online training and serving, and details of the first layer optimization for Pre-ranking models.
-Appendix B More Details about the Applications of Scaling Law
-B.1. ROI-Constrained Model Designing
-We set an upper limit on the model FLOPs to 143M and set the step size of the grid search as 16 and 128 for the feature embedding dimensions and the unit number of MLP layers, respectively.
+For both training and serving, we use T4 or A10 GPUs, with each GPU machine equipped with 128 CPU cores.
+Both the training and serving frameworks are developed based on TensorFlow [2] or PyTorch [37].
+Notably, the GPU models used for the experiments in Appendix C.1.2 differ from those used in standard online deployment.
+In Appendix C.1.2, to evaluate the performance of larger Transformer models, we utilize state-of-the-art, flagship-class GPUs rather than the T4 or A10 GPUs typically used in production.
+Due to company policies, certain detailed hardware specifications remain confidential and are not disclosed.
+Figure 11: The pipeline of online training and serving, and details of the first layer optimization for Pre-ranking models.
+D.2 An Detailed Illustration of MCET
+In this section, we provide a detailed illustration of the Machine Cost Estimation Tool (MCET), which is implemented based on Algorithm 1.
+The MCET is designed to estimate the computation cost of deploying machine learning models without the need for actual online deployment.
+This tool is particularly useful in scenarios where rapid iteration and cost-effective deployment are critical, such as in large-scale advertising systems.
+The core idea of MCET is to simulate the execution of models on a specified hardware configuration and measure their performance in terms of Queries Per Second (QPS) under a given response time limit.
+By leveraging these simulations, MCET can estimate the number of machines required to serve each model while meeting the desired performance constraints.
+This approach eliminates the need for costly and time-consuming online deployment, providing a reliable and efficient alternative for cost estimation.
+The algorithm begins by initializing a base model $M_{0}$ and simulating its execution on the specified hardware configuration $\mathcal{MH}$.
+The QPS of the base model $QPS_{0}$ is measured under the response time limit $T_{\text{limit}}$.
+Using this measurement, the total QPS capacity of the base model setup is calculated as $QPS_{\text{total}}=req_{0}\times QPS_{0}$, where $req_{0}$ is the number of machines allocated to the base model.
+Subsequently, the algorithm iterates over the remaining models, loading their meta files and simulating their execution on the same hardware configuration.
+For each model, the QPS $QPS_{i}$ is measured under the same response time limit, and the required number of machines $req_{i}$ is calculated as $req_{i}=\frac{QPS_{\text{total}}}{QPS_{i}}$.
+The results are stored in a list and returned as the final output.
+The MCET tool, which implements this algorithm, allows us to estimate the machine requirements for tens of models within a short time frame.
+For instance, using a single machine with a single GPU, we can estimate the required number of machines for a single TensorFlow meta-file within half an hour.
+This capability significantly reduces the time and cost associated with model deployment and iteration.
+Furthermore, by combining the estimated machine requirements with the unit price of the specified hardware configuration, we can accurately calculate the expected machine cost.
+This approach is applicable to both training and serving cost estimation, with the algorithm being executed using the meta files specific to each process.
+For training cost estimation, labels are randomly constructed to simulate the training environment.
+In summary, MCET provides a standardized, reproducible, and efficient method for estimating computation costs, enabling rapid and cost-effective model deployment in real-world applications.
+Algorithm 1 Cost Estimation by Offline Simulation Testing
+0: The number $n$ of models (denoted as $n$), and meta files of the models (denoted as $\{meta_{i}|0\leq i<n\}$).
+One specified machine $\mathcal{MH}$ for online serving.
+The machine number $req_{0}$ of the online base model $M_{0}$.
+The desired response time limit $T_{\text{limit}}$.
+0: The estimated machine number of each model for online serving, denoted as $\{req_{i}|1\leq i<n\}$.
+1: Initialize an empty list resultsresults to store the results.
+2: Simulate the execution of the base model $M_{0}$ using the specified hardware configuration $\mathcal{MH}$.
+3: Measure the QPS ($QPS_{0}$) of the base model $M_{0}$ under the given response time limit $T_{\text{limit}}$, with the input being randomly initialized, fixed-length tensors.
+4: Calculate the total QPS capacity of the base model setup: $QPS_{\text{total}}=req_{0}\times QPS_{0}$.
+5:for $i=1$ to n1n-1 do
+6:  Load the meta file $meta_{i}$ of the $i$-th model.
+7:  Simulate the model’s execution using the specified hardware configuration $\mathcal{MH}$.
+8:  Measure the QPS ($QPS_{i}$) of the $i$-th model under the given response time limit $T_{\text{limit}}$, with the input being randomly initialized, fixed-length tensors.
+9:  Calculate the required number of machines for the $i$-th model: $req_{i}=\frac{QPS_{\text{total}}}{QPS_{i}}$.
+10:  Append the result $req_{i}$ to the resultsresults list.
+11:end for
+12:return resultsresults
+D.3 ROI-Constrained Model Designing
+Regrading Case Study 1: We set an upper limit on the model FLOPs to 143M and set the step size of the grid search as 16 and 128 for the feature embedding dimensions and the unit number of MLP layers, respectively.
 This larger step size is chosen because smaller increments have minimal impact on the total FLOPs and result in negligible increases in expected revenue according to the scaling law.
-The feature embedding dimension ($emb\_dim$) in the grid search starts at 16.
-The input size of MLP models is determined by multiplying the $emb\_dim$ by the number of sparse features, which is 200 in our experiments, plus an additional 128-dimensional dense feature.
+The feature embedding dimension (emb_dimemb\_dim) in the grid search starts at 16.
+The input size of MLP models is determined by multiplying the emb_dimemb\_dim by the number of sparse features, which is 200 in our experiments, plus an additional 128-dimensional dense feature.
 For the MLP layers (excluding the output layer, which is fixed at 1), the number of units starts at 128, with the constraint that the number of units in any layer does not exceed the number of units in the previous layer.
 The size of the first layer outputs does not exceed 1024, because this is the storage limit of the embedding producer.
-Additionally, we enforce that the number of units in any layer (except the output layer) is at least 1/20 of the number of units in the previous layer as the reason discussed in appendix C.1.
+Additionally, we enforce that the number of units in any layer (except the output layer) is at least 1/20 of the number of units in the previous layer, as the reason discussed in appendix E.1.
 We used 10 T4-GPU machines to perform a grid search over approximately 1000 configurations and found an approximate optimal solution in about two days.
-B.2. Multi-scenario Resource Allocation
-We adopted an MLP model for the Pre-ranking stage and a twin-tower model for the Matching stage.
-Through offline experiments, we obtained the key functions (including $G$, $BNSL$, etc.) in the scaling laws for both the Matching and Pre-ranking stages using our proposed paradigm.
+Regrading Case Study 2: We observe from the fitted data points that models in the 0.05B to 0.1B parameter range offer favorable cost-performance trade-offs.
+Leveraging the scaling law for Transformer-based models, we therefore conduct a grid search within this interval with a step size of 0.01B.
+Ultimately, a 0.08B-parameter model was deployed.
+Given that the model size is moderate, we are able to use A10 GPUs for both training and serving.
+Notably, the scaling law we fit remains in terms of FLOPs and $R/R^{*}$.
+When the average number of tokens per query is known, there exists a straightforward relationship between model parameter count (in billions) and FLOPs.
+D.4 Multi-scenario Resource Allocation
+Regrading Case Study 3: As mentioned, we adopted an MLP model for the Pre-ranking stage and a DSSM model for the Matching stage.
+Through offline experiments, we obtained the key functions (including $G$, BNSLBNSL, etc.) in the scaling laws for both the Matching and Pre-ranking stages using our proposed paradigm.
 Due to the different model architectures, their machine costs vary in the same size.
-We used the machine cost estimation tool to additionally measure the machine costs of the twin-tower model at different sizes.
-Similarly to the estimation of the MLP model, we solved the estimated machine costs of approximately 1000 different sizes of twin-tower models based on the algorithm 1 using 10 T4-GPU machines for about 2 days.
-Appendix C Limitations and Future Work
-C.1. Optimal Model Designing with Given FLOPs
+We used the machine cost estimation tool to additionally measure the machine costs of the DSSM model at different sizes.
+Similarly to the estimation of the MLP model, we solved the estimated machine costs of approximately 1000 different sizes of DSSM models based on the algorithm 1 using 10 T4-GPU machines for about 2 days.
+Regrading Case Study 4: For $Scene_{1}$ and $Scene_{2}$, we perform proportional scaling of the base models in both the Pre-ranking and Matching stages.
+The scaling ratio ranges from 0.1 to 3.0 with a step size of 0.1.
+Based on the grid search results and the overall ROI across stages, we ultimately decide to reduce the model size in $Scene_{1}$ by 10% and increase the model size in $Scene_{2}$ by 50%.
+Appendix E Limitations and Future Work
+E.1 Optimal Model Designing with Given FLOPs
 One remaining question is how to give optimal unit distributions of layers under a given FLOPs.
@@
 More in-depth conclusions are left for future work and are beyond the scope of this paper.
-C.2. New Foundation Models and Joint Laws
-While the transformer has become a foundational model in the NLP domain, capable of being applied to almost all tasks and scenarios, the recommendation and advertising fields lack such a foundation model.
-This absence means that the scaling law parameters obtained in one scenario cannot be readily reused in another.
-In addition, our scaling laws show that the benefits of increasing FLOPs converge to zero much more rapidly.
-It’s different from the conclusion in NLP area, where increasing the FLOPs by an order of magnitude typically results in consistent performance improvements.
-An important reason might be that MLP models in online advertisement retrieval are not as effective foundation models as transformers in NLP.
-However, it is worth noting that transformers do not universally outperform MLPs in all advertising scenarios, especially when considering the ROI of the machine.
-Given these challenges, developing more advanced and universal foundation models with better scaling performance (e.g., recent works (Zhai et al., 2024; Zhang et al., 2024b; Guo et al., 2024)) is a crucial direction for future research in online advertisement.
-Additionally, obtaining joint laws considering multiple factors such as data, model size, and computational resources is significantly more costly in industrial advertising settings than obtaining single-variable scaling laws.
+E.2 Jow Laws Identification and Application
+Obtaining joint laws considering multiple factors, such as data, model size, and computational resources, is significantly more costly in industrial advertising settings than obtaining single-variable scaling laws.
 Developing a cost-effective method for deriving joint laws is therefore an important area for future investigation.
-The high cost of obtaining joint laws is also one of the motivations behind our exploration of unified foundation models.
+An especially promising direction for application is to model the joint law between model size and the computational quota allocation across stages in a cascade ranking system.
+Such a law could help answer critical questions: How should we rebalance ranking, pre-ranking, and matching stage budgets when scaling up the overall model?
+At what point does increasing model size in one stage yield diminishing returns due to bottlenecks in another?
+By formalizing these interactions, joint laws can deepen our understanding of cascade systems and guide principled, globally optimal configuration decisions with only a given budget.
+E.3 Beyond Compute Scaling
+Although our optimization framework has achieved an overall 5.1% improvement in ad revenue, a substantial portion of this gain arises not from breakthroughs in model architecture, but from correcting historically suboptimal resource allocation and coarse-grained manual tuning.
+Our results indicate that, unlike in NLP, simply scaling up computation alone is unlikely to bring similar levels of success to the advertising domain for models such as MLP, DSSM, or Transformer.
+Therefore, the next major leap in performance will likely require more disruptive innovations.
+One path is architectural rethinking—designing models with inherently better scaling efficiency, such as sparse, modular architectures [33].
+Another is a data-centric paradigm shift: leveraging large language models or product-level innovations to improve the quality, semantic richness, and diversity of training signals.
+In this view, "scaling" must evolve beyond model size to include the scaling of data intelligence.
+Only through such foundational changes can we hope to achieve the kind of sustained, huge progress seen in NLP domains.